### v1

# Lighter, Better, Faster Multi-Source Domain Adaptation with Gaussian Mixture Models and Optimal Transport

Abstract
In this paper, we tackle Multi-Source Domain Adaptation (MSDA), a task in transfer learning where one adapts multiple heterogeneous, labeled source probability measures towards a different, unlabeled target measure. We propose a novel framework for MSDA, based on Optimal Transport (OT) and Gaussian Mixture Models (GMMs). Our framework has two key advantages. First, OT between GMMs can be solved efficiently via linear programming. Second, it provides a convenient model for supervised learning, especially classification, as components in the GMM can be associated with existing classes. Based on the GMM-OT problem, we propose a novel technique for calculating barycenters of GMMs. Based on this novel algorithm, we propose two new strategies for MSDA: GMM-WBT and GMM-DaDiL. We empirically evaluate our proposed methods on four benchmarks in image classification and fault diagnosis, showing that we improve over the prior art while being faster and involving fewer parameters.
Keywords:
Domain Adaptation Optimal Transport Gaussian Mixture Models.
1 Introduction
Supervised learning models, especially deep neural nets, rely on large amounts of labeled data to learn a function that reliably predicts on unseen data. This property is known as generalization. However, these models are subject to performance degradation, when the conditions upon which test data is acquired changes. This issue is known in the literature as distributional, or dataset shift [22].
Under distributional shift, a possible solution is to acquire a new labeled dataset under the new conditions. This solution is, in many cases such as fault diagnosis [16], costly or infeasible. A different approach, known as Domain Adaptation (DA), consists of collecting an unlabeled target domain dataset, for which the knowledge in the source domain dataset is transferred to [19]. A way to further enhance this adaptation is to consider multiple related, but heterogeneous sources, which is known as Multi-Source DA (MSDA) [3].
In the context of DA, a prominent framework is Optimal Transport (OT) [30], which is a field of mathematics concerned with the displacement of mass at least effort. This theory has been applied for DA in multiple ways, especially by (i) mapping samples between domains [2] and (ii) learning invariant representations [25]. For MSDA, OT has been used for aggregating the multiple source domains into a barycentric domain [14, 13], which is later transported to the target domain, or by weighting source domain measures [27].
This work considers the problem of Wasserstein Dictionary Learning (WDL), which was initially proposed by [24] for histogram data. This problem was later generalized by [12], for empirical measures, which allowed its application to MSDA. In [12], one learns to express domains in MSDA as a barycenter of atom measures, which have a free, learnable support. As a result, the work of [12] learns how to interpolate distributional shift between the measures in MSDA.
However, previous algorithms relying on Wasserstein barycenters, such as Wasserstein Barycenter Transport (WBT) [14, 13] and Dataset Dictionary Learning (DaDiL) [12], are limited in scale, since the number of points the support of the empirical measures scale with the number of samples in the original datasets. As a consequence, previous works such as [14, 13] are limited to small scale datasets, or rely on mini-batch optimization [12], which introduces artifacts in the OT. To tackle the aforementioned limitations, in this paper we propose a novel, parametric framework for barycentric-based MSDA algorithms. based on recent developments in OT between Gaussian Mixture Models (GMMs) [5]. We present an overview of our methods in figure 1.
(a) GMM-WBT
(b) GMM-DaDiL
Figure 1: Overview of proposed methods. Squares represent actual domains, circles represent barycenters and triangles represent learned measures. Blue and orange elements represent labeled and unlabeled measures respectively. In GMM-WBT, a labeled GMM is determined for the target domain by transporting the barycenter of sources. In GMM-DaDiL, we learn to express each domain as a barycenter of learned GMMs, called atoms, through dictionary learning.
Summary of Contributions. Our contributions are threefold: 1. We propose a novel strategy for mapping the parameters of GMMs using OT (section 3.1 and theorem 3.1); 2. We propose a novel algorithm for computing mixture-Wasserstein barycenters of GMMs (algorithm 1 in section 3.3); 3. We propose an efficient parametric extension of the WBT and DaDiL algorithms based on GMMs (section 3.4). We highlight that, while GMMs were previously employed in single source DA [8], to the best of our knowledge this is the first work to leverage GMM-OT for MSDA.
Paper Organization. The rest of this paper is divided as follows. Section 2 covers the background behind our method. Section 3 covers our methodological contributions. Section 4 explores the empirical validation of our method with respect other OT-based MSDA algorithms, where we show that our methods significantly outperform prior art. Finally, section 5 concludes this paper.
2 Preliminaries
In this section, we cover the necessary background for our paper. We divide our discussion into Gaussian mixtures, domain adaptation, optimal transport and dataset dictionary learning.
Gaussian Mixtures. We denote the set of probability measures over a set $\mathcal{X}$ as $\mathbb{P}(\mathcal{X})$. A Gaussian measure corresponds to $P_{\theta}\in\mathbb{P}(\mathcal{X})$ with the following density,
$f_{\theta}(\mathbf{x})$ $=\dfrac{1}{\sqrt{(2\pi)^{d}\det(\mathbf{C}^{(P)})}}\exp\biggr{(}-\dfrac{1}{2}(\mathbf{x}-\mathbf{m}^{(P)})^{T}(\mathbf{C}^{(P)})^{-1}(\mathbf{x}-\mathbf{m}^{(P)})\biggr{)},$
where $\theta=(\mathbf{m}^{(P)},\mathbf{C}^{(P)})$ are the mean vector $\mathbf{m}^{(P)}\in\mathbb{R}^{d}$ and the covariance matrix $\mathbf{C}^{(P)}\in\mathbb{S}^{d}=\{\mathbf{C}\in\mathbb{R}^{d\times d}:\mathbf{C}=\mathbf{C}^{T}\text{ and }\mathbf{x}\mathbf{C}\mathbf{x}^{T}>0,\forall\mathbf{x}\in\mathbb{R}^{d}\setminus\{\mathbf{0}\}\}$. We generally denote $P_{\theta}=\mathcal{N}(\mathbf{m}^{(P)},\mathbf{C}^{(P)})$. In addition, let $K\geq 1$ be an integer. A GMM over $\mathbb{R}^{d}$ is a probability measure $P_{\theta}\in\mathbb{P}(\mathbb{R}^{d})$ such that,
$P_{\theta}=\sum_{k=1}^{K}p_{k}P_{k}\text{, where }P_{k}=\mathcal{N}(\mathbf{m}_{k}^{(P)},\mathbf{C}_{k}^{(P)})\text{, and }\mathbf{p}\in\Delta_{K},$ (1)
where $\Delta_{K}=\{\mathbf{p}\in\mathbb{R}^{K}:\sum_{k=1}^{K}p_{k}=1\}$. Following [5], we denote the subset of $\mathbb{P}(\mathbb{R}^{d})$ of probability measures which can be written as Gaussian mixtures with less than $K$ components by $\text{GMM}_{d}(K)$, and $\text{GMM}_{d}(\infty)=\cup_{k\geq 0}\text{GMM}_{d}(K)$.
Given data points $\{\mathbf{x}_{i}^{(P)}\}_{i=1}^{n}$ i.i.d. from $P$, one can determine the parameters $\theta$ through maximum likelihood,
$\theta^{\star}$ $=\underset{\theta\in\Theta}{\text{argmax}}\,\sum_{i=1}^{n}\log P_{\theta}(\mathbf{x}_{i}^{(P)}),$ (2)
where $\Theta=\{\{p_{k},\mathbf{m}_{k}^{(P)},\mathbf{C}_{k}^{(P)}\}_{k=1}^{K}:\sum_{k}p_{k}=1\text{ and }\mathbf{C}_{k}^{(P)}\in\mathbb{S}^{d}\}$. While equation 2 has no closed-form solution, one can solve this optimization problem through the celebrated Expectation-Maximization (EM) algorithm [6].
Learning Theory and Domain Adaptation. In this paper, we focus on the problem of classification. Given a feature space $\mathcal{X}=\mathbb{R}^{d}$ and a label space $\mathcal{Y}=\{1,\cdots,n_{cl}\}$, this problem corresponds to finding $h\in\mathcal{H}\subset\mathcal{X}^{\mathcal{Y}}$ that correctly classifies data $\{(\mathbf{x}_{i}^{(Q)},y_{i}^{(Q)})\}_{i=1}^{n}$.
In this paper, we use the Empirical Risk Minimization (ERM) framework [28], as it is useful for domain adaptation theory. As follows, one assumes $\mathbf{x}_{i}^{(Q)}\overset{iid}{\sim}Q$, for a measure $Q\in\mathbb{P}(\mathcal{X})$, and $h_{0}\in\mathcal{X}^{\mathcal{Y}}$ such that $y_{i}^{(Q)}=h_{0}(\mathbf{x}_{i}^{(Q)})$. $h_{0}$ is called ground-truth labeling function. Given a loss function $\mathcal{L}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, a classifier may be defined through risk minimization, i.e., $h^{\star}=\text{argmin}_{h\in\mathcal{H}}\mathcal{R}_{Q}(h)$, for $\mathcal{R}_{Q}(h)=\mathbb{E}_{Q}[\mathcal{L}(h(\mathbf{x}),h_{0}(\mathbf{x}))]$. This strategy is oftentimes impractical as $Q$ and $h_{0}$ are unknown. As a result, one resorts to the minimization of the empirical risk, i.e., $\hat{h}=\text{argmin}_{h\in\mathcal{H}}\hat{\mathcal{R}}_{Q}(h)$, where ^$\hat{\mathcal{R}}_{Q}(h)=\dfrac{1}{n}\sum_{i=1}^{n}\mathcal{L}(h(\mathbf{x}_{i}^{(Q)}),y_{i}^{(Q)})$.
From a theoretical standpoint, this framework is useful because $\mathcal{R}_{Q}$ is bounded by ^$\hat{\mathcal{R}}_{Q}$ and a complexity term depending on the number of samples $n$, and the Vapnik-Chervonenkis dimension of $\mathcal{H}$ [28, Section 6]. As a result, $\hat{h}$ minimizing the empirical risk is guaranteed to generalize to unseen samples of $P$.
Nevertheless, the assumption that unseen examples come from a fixed measure $P$ is seldom verified in practice [22], since the conditions upon which data is acquired may change. In this case, models are required to adapt to new data, but at the same time re-training a model from the scratch is likely costly and data intensive. A solution consists of using transfer learning [19], in which one re-uses knowledge from a source domain or task to facilitate the learning on a target domain or task.
In transfer learning, a domain is a pair $(\mathcal{X},Q(X))$ of a feature space and a (marginal) probability measure. Likewise, a task is a pair $(\mathcal{Y},Q(Y|X))$ of a label space and a conditional probability measure. Domain adaptation is a case in which one has two domains $(\mathcal{X},Q_{S}(X))$, $(\mathcal{X},Q_{T}(X))$, a single task $(\mathcal{Y},Q(Y|X))$, and $Q_{S}(X)\neq Q_{T}(X)$. Furthermore, multi-source domain adaptation supposes multiple source domain measures, i.e., $Q_{S_{1}},\cdots,Q_{S_{N}}$, with $Q_{S_{i}}\neq Q_{S_{j}}$, and $Q_{S_{i}}\neq Q_{T}$. To reflect the idea that acquiring new data is costly, we have an unsupervised scenario. In this case, we have $N$ labeled source datasets $\{(\mathbf{x}_{i}^{(Q_{S_{\ell}})},y_{i}^{(Q_{S_{\ell}})})\}_{i=1}^{n_{\ell}}$, and an unlabeled target dataset $\{\mathbf{x}_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$. Our goal is to learn a classifier on $Q_{T}$ by leveraging the knowledge from the source domains.
Optimal Transport is a field of mathematics concerned with the displacement of mass at least effort [21, 15]. Given probability measures $P,Q\in\mathbb{P}(\mathcal{X})$, the Monge formulation [21, Section 2.2.] of OT seeks for a mapping $T$,
[…]
where $T_{\sharp}$ is the push-forward mapping of $T$, i.e., $T_{\sharp}P(A)=P(T^{-1}(A))$. Nonetheless, this problem poses technical difficulties, mainly due the constraint $T_{\sharp}P=Q$. A more tractable formulation was proposed by Kantorovich [21, Section 2.3.], and relies on OT plans,
[…]
where $\Gamma(P,Q)=\{\gamma\in\mathbb{P}(\mathcal{X}\times\mathcal{X}):\int_{\mathcal{X}}\gamma(A,z)=P(A)\text{, and }\int_{\mathcal{X}}\gamma(x,B)=Q(B)\}$ is called the transportation polytope. There is a metric between probability measures, associated with OT, called Wasserstein distance [30]. As such, let $c(x,z)=d(x,z)^{\alpha}$ for $\alpha\in[1,\infty)$, where $d$ is a metric on $\mathcal{X}$, then,
$\mathcal{W}_{c,\alpha}(P,Q)=\biggr{(}\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\int_{\mathcal{X}}\int_{\mathcal{X}}c(x,z)d\gamma(x,z)\biggr{)}^{\nicefrac{{1}}{{\alpha}}}.$ (5)
When $\mathcal{X}=\mathbb{R}^{d}$, a common choice is $c(\mathbf{x},\mathbf{z})=\lVert\mathbf{x}-\mathbf{z}\rVert_{2}^{\alpha}$, for which we omit the subscript $c$. Furthermore, common values for $\alpha$ include $1$ and $2$. Throughout this paper we adopt the Euclidean metric and $\alpha=2$.
While equation 4 is hard to solve for general $P$ and $Q$, it has closed-form solution for Gaussian measures [26]. As such, let $P=\mathcal{N}(\mathbf{m}^{(P)},\mathbf{C}^{(P)})$ (resp. $Q$). Under these conditions, for $\mathbf{C}^{(P)}=\mathbf{S}^{(P)}(\mathbf{S}^{(P)})^{T}$,
$\mathcal{W}_{2}^{2}(P,Q)=\lVert\mathbf{m}^{(P)}-\mathbf{m}^{(Q)}\rVert^{2}_{2}+\text{Tr}\biggr{(}\mathbf{C}^{(P)}+\mathbf{C}^{(Q)}-2(\mathbf{S}^{(P)}\mathbf{C}^{(Q)}\mathbf{S}^{(P)})^{\nicefrac{{1}}{{2}}}\biggr{)},$
This expression can be further simplified for axis-aligned Gaussians, i.e., $\mathbf{S}^{(P)}=\text{diag}(\mathbf{s}^{(P)})$, with $\mathbf{s}^{(P)}\in\mathbb{R}^{d}_{+}$,
$\mathcal{W}_{2}^{2}(P,Q)$ $=\lVert\mathbf{m}^{(P)}-\mathbf{m}^{(Q)}\rVert^{2}_{2}+\lVert\mathbf{s}^{(P)}-\mathbf{s}^{(Q)}\rVert_{2}^{2}.$ (6)
Henceforth, we assume axis-aligned Gaussian measures.
In this paper, we use the GMM-OT framework of [5], which is convenient to our setting for 2 reasons. First, they are able to represent measures with sub-populations, such as those commonly encountered in classification and domain adaptation. Second, they yield a tractable OT problem, when $\gamma$ is further restricted to be a GMM itself, that is,
$\omega^{\star}$ $=\text{GMMOT}(P,Q)=\underset{\omega\in\Gamma(\mathbf{p},\mathbf{q})}{\text{argmin}}\,\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}\mathcal{W}_{2}(P_{i},Q_{j})^{2},$ (7)
where the OT plan is given by $\gamma^{\star}=\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}^{\star}f_{\theta}(\mathbf{x})\delta(\mathbf{y}-T_{ij}(\mathbf{x}))$ and $\mathcal{W}_{2}(P_{i},Q_{j})^{2}$ is the Wasserstein distance between components $P_{i}$ and $Q_{j}$ (c.f., eq. 6). Furthermore, the GMMOT problem defines the Mixture-Wasserstein distance [5],
$\mathcal{MW}_{2}(P,Q)^{2}$ $=\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathcal{W}_{2}(P_{i},Q_{j})^{2}.$ (8)
Dataset Dictionary Learning [12] is an OT-based framework for expressing probability measures as barycenters of synthetic measures, called atoms. The framework is inspired by dictionary learning literature [24]. The authors introduce atoms $\mathcal{P}=\{\hat{P}_{c}\}_{c=1}^{C}$ and barycentric coordinates $\Lambda=\{\lambda_{\ell}\}_{\ell=1}^{N}$, such that each measure in MSDA is expressed as a Wasserstein barycenter $\hat{Q}_{\ell}=\mathcal{B}(\lambda_{\ell};\mathcal{P})$. This framework leads to the following optimization problem,
$\mathcal{L}(\Lambda,\mathcal{P})$ $=\underset{\Lambda,\mathcal{P}}{\text{argmin}}\,\mathcal{W}_{2}(\hat{Q}_{T},\mathcal{B}(\lambda_{\ell};\mathcal{P}))+\sum_{\ell=1}^{N}\mathcal{W}_{c,2}(\hat{Q}_{\ell},\mathcal{B}(\lambda_{\ell};\mathcal{P}))^{2},$ (9)
where $c\biggr{(}(\mathbf{x}^{(Q_{\ell})},\mathbf{y}^{(Q_{\ell})}),(\mathbf{x}^{(B_{\ell})},\mathbf{y}^{(B_{\ell})})\biggr{)}=\lVert\mathbf{x}^{(Q_{\ell})}-\mathbf{x}^{(B_{\ell})}\rVert_{2}^{2}+\beta\lVert\mathbf{y}^{(Q_{\ell})}-\mathbf{y}^{(B_{\ell})}\rVert_{2}^{2}$, and $\beta\geq 0$ is a constant expressing how costly it is to move samples from different classes. This framework makes it easier to express the distributional shift between the different $\mathcal{Q}=\{\hat{Q}_{1},\cdots,\hat{Q}_{N_{S}},\hat{Q}_{T}\}$. Especially, since $\hat{B}_{\ell}=\mathcal{B}(\lambda_{\ell};\mathcal{P})$ is labeled, one can synthesize labeled target domain data by reconstructing the target measure with $\lambda_{T}:=\lambda_{N+1}$.
3 Methodological Contributions
3.1 First Order Analysis of $\mathcal{MW}_{2}$
In this section we analyze $P\mapsto\mathcal{MW}_{2}(P,Q)$, for a fixed $Q$. We are particularly interested on how to map the components of $P$ towards $Q$, while minimizing this distance. The following theorem provides us a strategy, [^1] We refer readers to our supplementary materials for all proofs and derivations.
Theorem 3.1
Let $P$ and $Q$ be two GMMs with components $P_{i}=\mathcal{N}(\mathbf{m}_{i}^{(P)},(\mathbf{s}_{i}^{(P)})^{2})$ (resp. $Q_{j}$) and $\omega^{\star}$ be the solution of eq. 8. The first-order optimality conditions of $\mathcal{MW}_{2}^{2}$, with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by,
$\mathbf{m}_{i}=T_{\omega^{\star}}(\mathbf{m}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{m}_{j}^{(Q)}\text{, and }\mathbf{s}_{i}=T_{\omega^{\star}}(\mathbf{s}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{s}_{j}^{(Q)},$ (10)
where $\omega^{\star}$ is the solution of eq. 7.
Equation 10 is similar to the barycentric mapping in Empirical Optimal Transport (EOT) [2, eq. 13], which serves as an approximation for the Monge mapping between $P$ and $Q$. In our case, the barycentric mappings act on the parameters of the GMM, rather than on its samples. Theorem 3.1 will be useful in the calculation of $\mathcal{MW}_{2}$ barycenters.
3.2 Supervised Mixture-Wasserstein Distances
In this paper, we consider supervised learning problems. As such, it is necessary to equip the components of GMMs with labels that represent the classes in the datasets. We propose doing so through a simple heuristic, based on the law of total probability,
$P(y|k)$ $=\sum_{i=1}^{n}P_{\theta}(\mathbf{x}_{i}|k)P(y|\mathbf{x}_{i})=\sum_{i=1}^{n}\biggr{(}\dfrac{p_{k}P_{k}(\mathbf{x}_{i})}{\sum_{k^{\prime}}p_{k^{\prime}}P_{k^{\prime}}(\mathbf{x}_{i})}\biggr{)}P(y|\mathbf{x}_{i}),$ (11)
where $P_{\theta}(\mathbf{x}_{i}|k)$ corresponds to the probability that $\mathbf{x}_{i}$ comes form the $k-$th component of the GMM. Henceforth, we denote the soft-labels of components by $\mathbf{y}_{i}^{(P)}$. We may use the labeled GMM for obtaining a classifier, by sampling $\mathbf{x}\sim P_{\theta}$. If $\mathbf{x}$ is sampled from the $k-$th component, we simply assign $P(y|k)$ as its soft-label. We illustrate our strategy in figure 2.
(a) Gaussian Mixture $P_{\theta}$
(b) $\text{argmax}_{k}P_{\theta}(\mathbf{x}|k)$.
(c) Labeled GMM
Figure 2: Illustration of the labeling process of Gaussian mixtures. (a) First, the mixture $P_{\theta}$ is learned. (b) Then, one estimates $P_{\theta}(\mathbf{x}|k)$ for each $\mathbf{x}_{i}$. (c) The label of each component can be estimated using the law of total probability.
Similarly to EOT, when the mixtures $P$ and $Q$ are labeled, one needs to take into account the labels in the ground-cost. Given $\beta>0$, we propose the following distance between labeled GMMs,
$\mathcal{SMW}_{2}(P,Q)^{2}$ $=\underset{\omega\in\Gamma(\mathbf{p},\mathbf{q})}{\text{argmin}}\,\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}(\mathcal{W}_{2}(P_{i},Q_{j})^{2}+\beta\lVert\mathbf{y}_{i}^{(P)}-\mathbf{y}_{j}^{(Q)}\rVert_{2}^{2}).$ (12)
While simple, using an Euclidean distance for the soft-labels allows us to derive similar first-order conditions for $\mathcal{SMW}_{2}$,
Theorem 3.2
Under the same conditions of theorem 3.1, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{y}_{i}^{(P)}$ and $\mathbf{y}_{j}^{(Q)}$. The first order optimality conditions of $\mathcal{SMW}_{2}$ with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by eq. 10. Furthermore, for $\mathbf{y}_{i}$,
$\hat{\mathbf{y}}_{i}=T_{\omega}(\mathbf{y}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{y}_{j}^{(Q)}.$ (13)
3.3 Mixture Wasserstein Barycenters
In this section, we detail our first methodological contribution: a new algorithm for computing barycenters of GMMs under the $\mathcal{MW}_{2}$ and $\mathcal{SMW}_{2}$ metrics. As such, we adapt the definition of [1],
Definition 1
Given $C\geq 1$ GMMs $\mathcal{P}=\{P_{c}\}_{c=1}^{C}$ and a vector of barycentric coordinates $\lambda\in\Delta_{C}$, the $\mathcal{SMW}_{2}$ barycenter is given by,
$P^{\star}=\mathcal{B}(\lambda,\mathcal{P})=\underset{B\in\text{GMM}_{d}(\infty)}{\text{argmin}}\,\biggr{\{}\mathcal{L}(B)=\sum_{c=1}^{C}\lambda_{c}\mathcal{SMW}_{2}(B,P_{c})^{2}\biggr{\}}.$ (14)
When the GMMs in $\mathcal{P}$ are unlabeled, one may define, by analogy, a barycenter under the $\mathcal{MW}_{2}$. Henceforth we describe an algorithm for labeled GMMs, but its extension for unlabeled GMMs is straightforward.
Inspired by previous results in empirical Wasserstein barycenters [4, 12], we propose a novel strategy for computing $\mathcal{B}(\lambda,\mathcal{P})$. Our idea consists on initializing the parameters $\{(\mathbf{m}_{i}^{(B)},\mathbf{s}_{i}^{(B)},\mathbf{y}_{i}^{(B)})\}_{i=1}^{K_{B}}$ at random. We then calculate $C$ GMM-OT transport plans $\{\omega^{(c)}\}_{c=1}^{C}$ between $P_{c}$ and $B$. Next, we use equations 10 and 13 to update the parameters of the barycenter $B$. We refer readers to our supplementary materials for further details. This strategy effectively minimizes $\mathcal{L}$ in eq. 14. We show our strategy in Algorithm 1.
1 function mixture_wasserstein_barycenter($\{\mathbf{M}^{(P_{c})}\}_{c=1}^{C}$, $\{\mathbf{S}^{(P_{c})}\}_{c=1}^{C}$, $\{\mathbf{Y}^{(P_{c})}\}_{c=1}^{C}$, $\{\lambda_{c}\}_{c=1}^{C}$, $\tau$, $N_{it}$)
2       $\mathbf{m}_{i}^{(B)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, $\mathbf{s}_{i}^{(B)}:=1$ and $\mathbf{y}_{i}^{(B)}:=\nicefrac{{1}}{{n_{c}}}$
3       while $|L_{it}-L_{it-1}|\geq\tau$ and $it\leq N_{it}$ do
             // Compute GMM-OT plans
4             for $c=1,\cdots,C$ do
5                   $\omega^{(c,it)}=\text{GMMOT}(B,P_{c})$
6
            // Note: $\mathcal{W}_{2}(B_{i},P_{c,j})^{2}=\lVert\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})}\rVert_{2}^{2}+\lVert\mathbf{s}_{i}^{(B)}-\mathbf{s}_{j}^{(P_{c})}\rVert_{2}^{2}$
7             $L_{it}=\sum_{c=1}^{C}\lambda_{c}\sum_{i=1}^{K_{B}}\sum_{j=1}^{K_{P}}\omega^{(c,it)}_{ij}\biggr{(}(\mathcal{W}_{2}(B_{i},P_{c,j})^{2}+\lVert\mathbf{y}^{(B)}_{i}-\mathbf{y}^{(P_{c})}_{j}\rVert_{2}^{2})\biggr{)}$
             // Update barycenter parameters
8             $\mathbf{m}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{k}T_{\omega^{(k,it)}}(\mathbf{m}^{(B)}_{i})$
9             $\mathbf{s}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{k}T_{\omega^{(k,it)}}(\mathbf{s}^{(B)}_{i})$
10             $\mathbf{y}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{k}T_{\omega^{(k,it)}}(\mathbf{y}^{(B)}_{i})$
11
12      return $\mathbf{M}^{(B)}$, $\mathbf{S}^{(B)}$, $\mathbf{Y}^{(B)}$
13
Algorithm 1 $\mathcal{SMW}_{2}$ Barycenter of GMMs
3.4 Multi-Source Domain Adaptation through GMM-OT
In this section, we detail two contributions for MSDA based on GMM-OT: GMM-WBT and GMM-DaDiL. In both cases, we suppose access to $N$ labeled source measures $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N}$ and an unlabeled target measure $Q_{T}$. Contrary to the empirical versions of these algorithms [14, 13, 12], we assume that an axis-aligned GMM has been learned for each domain, including the target.
GMM-WBT. The intuition of this algorithm is transforming the MSDA scenario into a single-source one, by first calculating a Wasserstein barycenter of $B=\mathcal{B}(\mathds{1}_{N}/N;\mathcal{Q}_{S})$. After this step, WBT solves a single-source problem between $B$ and $Q_{T}$. When each $Q_{S_{\ell}}$ is a GMM, the parameters of $B$ are estimated through algorithm 1. Next, one solves for $\omega^{(T)}=\text{GMMOT}(B,Q_{T})$, so that the parameters of $B$ are transported towards $Q_{T}$ using theorems 3.1 and 3.2,
$\hat{\mathbf{m}}_{i}^{(Q_{T})}=K_{B}\sum_{j=1}^{K_{T}}\omega_{ij}^{(T)}\mathbf{m}_{j}^{(Q_{T})}\text{, and, }\hat{\mathbf{s}}_{i}^{(Q_{T})}=K_{B}\sum_{j=1}^{K_{T}}\omega_{ij}^{(T)}\mathbf{s}_{j}^{(Q_{T})}.$ (15)
With a labeled GMM, $\{\hat{\mathbf{m}}_{i}^{(Q_{T})},\hat{\mathbf{s}}_{i}^{(Q_{T})},\mathbf{y}_{i}^{(B)}\}_{i=1}^{K_{B}}$, on the target domain, we can learn a classifier on the target domain as explained in section 3.2.
1 function gmm_dadil($\{\mathbf{M}^{(Q_{S_{\ell}})}\}_{\ell=1}^{N}$, $\{\mathbf{S}^{(Q_{S_{\ell}})}\}_{\ell=1}^{N}$, $\{\mathbf{Y}^{(Q_{S_{\ell}})}\}_{\ell=1}^{N}$, $N_{it}$, $\eta$)
       // Initialization.
2       $\mathbf{m}_{i}^{(P_{k})}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, $\mathbf{s}_{i}^{(P_{k})}:=1$, $\mathbf{u}_{i}^{(P_{k})}:=\nicefrac{{1}}{{n_{c}}}$, and $\lambda_{\ell}=\nicefrac{{1}}{{K}}$
3       for $it=1,\cdots,N_{it}$ do
4             $L=0$
             // Evaluate supervised loss on sources
5             for $\ell=1,\cdots,N$ do
6                   $L\leftarrow L+\mathcal{SMW}_{2}(Q_{S_{\ell}},\mathcal{B}(\lambda_{\ell},\mathcal{P}))^{2}$
7
            // Evaluate unsupervised loss on targets
8             $L\leftarrow L+\mathcal{MW}_{2}(Q_{T},\mathcal{B}(\lambda_{T},\mathcal{P}))^{2}$
             // Gradient step
9             $\mathbf{x}_{j}^{(P_{k})}\leftarrow\mathbf{x}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{x}^{(P_{k})}_{j}}}$
10             $\mathbf{p}_{j}^{(P_{k})}\leftarrow\mathbf{u}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{u}^{(P_{k})}_{j}}}$
             // Note: we project variables $\mathbf{s}$ and $\lambda$.
11             $\mathbf{s}_{j}^{(P_{k})}\leftarrow\text{proj}_{\mathbb{R}^{d}_{+}}(\mathbf{s}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{s}^{(P_{k})}_{j}}})$
12             $\lambda_{\ell}\leftarrow\text{proj}_{\Delta_{K}}(\lambda_{\ell}-\eta\nicefrac{{\partial L}}{{\partial\lambda_{\ell}}})$
13
14      return $\mathcal{P},\Lambda$
15
Algorithm 2 GMM-Dataset Dictionary Learning
GMM-DaDiL. Our second algorithm consists of a parametric version for the DaDiL algorithm of [12]. The idea is to replace the atoms in $\mathcal{P}=\{\hat{P}_{c}\}_{c=1}^{C}$ by GMMs parametrized through $\Theta_{P}=\{(\mathbf{M}^{(P_{c})},\mathbf{S}^{(P_{c})},\mathbf{Y}^{(P_{c})})\}_{c=1}^{C}$. Learning a dictionary is thus equivalent to estimating these parameters, that is,
$\mathcal{L}(\Lambda,\Theta_{P})$ $=\underset{\Lambda,\mathcal{P}}{\text{argmin}}\,\mathcal{MW}_{2}(Q_{T},\mathcal{B}(\lambda_{T};\mathcal{P}))^{2}+\sum_{\ell=1}^{N}\mathcal{SMW}_{2}(Q_{\ell},\mathcal{B}(\lambda_{\ell};\mathcal{P}))^{2}.$ (16)
An important feature of this strategy is that the barycenters $\mathcal{B}$ inherit the labels from the atoms $\mathcal{P}$. As a result, we can estimate a labeled GMM for $Q_{T}$ in one step through $\mathcal{B}(\lambda_{T};\mathcal{P})$.
While eq. 16 does not have a closed-form solution, we optimize it through gradient descent. An advantage of the GMM modeling is that this optimization problem involves far less variables than DaDiL, hence we do not resort to mini-batches. We detail our strategy in Algorithm 2. Note that we need to enforce 3 kinds of constraints: (i) $\mathbf{s}_{i}^{(P_{c})}\in\mathbb{R}^{d}_{+}$, (ii) $\lambda_{\ell}\in\Delta_{C}$ and (iii) $\mathbf{y}_{i}^{(P_{c})}\in\Delta_{n_{cl}}$. For (i) and (ii), we use orthogonal projections into +$\mathbb{R}^{d}_{+}$ and $\Delta_{C}$ respectively. We additionally set $\mathbf{s}_{i}^{(P_{c})}\geq s_{min}$ for numerical stability. For (iii), we perform a change of variables $\mathbf{y}_{i}^{(P_{c})}=\text{softmax}(\mathbf{u}_{i}^{(P_{c})})$.
Once the dictionary $(\Lambda,\mathcal{P})$ is learned, we are able to reconstruct the domains in MSDA via the barycenter $\mathcal{B}(\lambda;\mathcal{P})$. We are especially interested in the target reconstruction $\lambda_{T}$, i.e., $\mathcal{B}(\lambda_{T};\mathcal{P})$. This barycenter is a labeled GMM (as we show in figure 5b). As a result, we can obtain labeled samples from this GMM, then use them to train a classifier that works on the target domain.
Faster Dataset Dictionary Learning. The computational complexity of an optimization step of algorithm 2 corresponds to $\mathcal{O}(N\times N_{it}\times C\times K^{3}\log K)$, i.e., we calculate $N$ barycenters of $C$ atoms. One should compare this complexity with that of DaDiL, i.e., $\mathcal{O}(N\times N_{it}\times M\times C\times n_{b}^{3}\log n_{b})$, where $M=\lceil\nicefrac{{n}}{{n_{b}}}\rceil$ is the number of mini-batches sampled at each iteration. In our experiments in section 4, we show that we achieve state-of-the-art performance with $K$ on the same order of magnitude as $n_{b}$ (e.g., a few hundred Gaussian components). As a result, we achieve a speed-up on the order of $M$ while solving an exact OT problem.
4 Experiments
In this section, we present our experiments. First, we present a toy example, where the source and target measures were generated by transforming an initial measure. As such, this toy example illustrates our goal to model distributional shift. Second, we experiment with benchmarks in MSDA.
Figure 3: Data used in the toy experiment. Each of these datasets was generated by applying an affine transformation to an initial dataset.
Toy Example. In figure 3 we show the generated datasets. Starting with GMM-WBT, figure 4(a) shows the learned GMMs for each dataset, in which the target GMM is not labeled. The barycenter of $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N}$ is shown in figure 4(b) (bottom-left). This barycenter is labeled. As a result, we may transfer its parameters to the target domain through GMM-OT (upper part), which leads to a labeled GMM in the target domain.
(a) GMMs learned on each domain.
(b) GMM-WBT.
Figure 4: Summary of GMM-WBT, in which a labeled GMM is determined on the target domain by transporting the parameters of the barycenter of sources.
Next, we show in figure 5 a summary for the GMM-DaDiL optimization process (figure 5(a)), and the reconstruction of target domain GMMs (figure 5(b)). Note that, as the training progresses, the reconstruction error and the negative log-likelihood of the GMMs decrease. As a result, GMM-DaDiL produces accurate, labeled GMMs for each domain. We provide further examples on the GMMs-DaDiL optimization in the supplementary materials. Next, we present our results on MSDA benchmarks.
(a) Optimization summary.
(b) Reconstructions ($it=200$).
Figure 5: Optimization of GMM-DaDiL and reconstruction of domain GMMs.
Experimental Setting. We compare our method to prior art. We focus on OT-based methods, such as WJDOT [27], WBT [14, 13] and DaDiL [12]. For completeness, we include recent strategies on deep MSDA that update the encoder network during the adaptation process, rather than using pre-extracted features. These are M3SDA [20], LtC-MSDA [31], KD3A [7] and Co-MDA [10]. We establish our comparison on 4 benchmarks, divided between visual domain adaptation (Office31 [23], Office-Home [29]) and cross-domain fault diagnosis (TEP [17] and CWRU). See table 1 for further details.
Table 1: Overview of benchmarks used in our experiments.
Benchmark Backbone Problem # Samples # Domains # Classes # Features
Office 31 ResNet50 Object Recognition 3287 3 31 2048
Office-Home ResNet101 Object Recognition 15500 4 65 2048
TEP CNN Fault Diagnosis 17289 6 29 128
CWRU MLP Fault Diagnosis 24000 3 10 256
As with previous works on OT-based MSDA, we perform domain adaptation on pre-extracted features. As such, we pre-train a neural network (called backbone) on the concatenation of source domain data, then we use it to extract the features from each domain. For visual adaptation tasks, we use ResNets [9], while for fault diagnosis, we use a CNN and a multi-layer perceptron, as in [12, 17]. We summarize our results in table 2(d).
[…]
Algorithm Ar Cl Pr Rw Avg. $\uparrow$
ResNet101 72.90 62.20 83.70 85.00 75.95
M3SDA 71.13 61.41 80.18 80.64 73.34
LtC-MSDA 74.52 60.56 85.52 83.63 76.05
KD3A 73.80 63.10 84.30 83.50 76.17
Co-MDA 74.40 64.00 85.30 83.90 76.90
WJDOT 74.28 63.80 83.78 84.52 76.59
WBT 75.72 63.80 84.23 84.63 77.09
DaDiL-E 77.16 64.95 85.47 84.97 78.14
DaDiL-R 75.92 64.83 85.36 85.32 77.86
GMM-WBT 75.31 64.26 86.71 85.21 77.87
GMM-DaDiL 77.16 66.21 86.15 85.32 78.81
(a) Office-Home.
Algorithm A D W Avg. $\uparrow$
ResNet50 67.50 95.00 96.83 86.40
M3SDA 66.75 97.00 96.83 86.86
LtC-MSDA 66.82 100.00 97.12 87.98
KD3A 65.20 100.0 98.70 87.96
Co-MDA 64.80 99.83 98.70 87.83
WJDOT 67.77 97.32 95.32 86.80
WBT 67.94 98.21 97.66 87.93
DaDiL-E 70.55 100.00 98.83 89.79
DaDiL-R 70.90 100.00 98.83 89.91
GMM-WBT 70.13 99.11 96.49 88.54
GMM-DaDiL 72.47 100.0 99.41 90.63
(b) Office 31.
Algorithm A B C Avg. $\uparrow$
MLP 70.90 $\pm$ 0.40 79.76 $\pm$ 0.11 72.26 $\pm$ 0.23 74.31
M3SDA 56.86 $\pm$ 7.31 69.81 $\pm$ 0.36 61.06 $\pm$ 6.35 62.57
LTC-MSDA 82.21 $\pm$ 8.03 75.33 $\pm$ 5.91 81.04 $\pm$ 5.45 79.52
KD3A 81.02 $\pm$ 2.92 78.04 $\pm$ 4.05 74.64 $\pm$ 5.65 77.90
Co-MDA 62.66 $\pm$ 0.96 55.78 $\pm$ 0.85 76.35 $\pm$ 0.79 64.93
WJDOT 99.96 $\pm$ 0.02 98.86 $\pm$ 0.55 100.0 $\pm$ 0.00 99.60
WBT 99.28 $\pm$ 0.18 79.91 $\pm$ 0.04 97.71 $\pm$ 0.76 92.30
DaDiL-R 99.86 $\pm$ 0.21 99.85 $\pm$ 0.08 100.00 $\pm$ 0.00 99.90
DaDiL-E 93.71 $\pm$ 6.50 83.63 $\pm$ 4.98 99.97 $\pm$ 0.05 92.33
GMM-WBT 100.00 $\pm$ 0.00 99.95 $\pm$ 0.07 100.00 $\pm$ 0.00 99.98
GMM-DaDiL 100.00 $\pm$ 0.00 99.95 $\pm$ 0.04 100.00 $\pm$ 0.00 99.98
(c) CWRU.
Algorithm Mode 1 Mode 2 Mode 3 Mode 4 Mode 5 Mode 6 Avg. $\uparrow$
CNN 80.82 $\pm$ 0.96 63.69 $\pm$ 1.71 87.47 $\pm$ 0.99 79.96 $\pm$ 1.07 74.44 $\pm$ 1.52 84.53 $\pm$ 1.12 78.48
M3SDA 81.17 $\pm$ 2.00 61.61 $\pm$ 2.71 79.99 $\pm$ 2.71 79.12 $\pm$ 2.41 75.16 $\pm$ 3.01 78.91 $\pm$ 3.24 75.99
KD3A 72.52 $\pm$ 3.04 18.96 $\pm$ 4.54 81.02 $\pm$ 2.40 74.42 $\pm$ 1.60 67.18 $\pm$ 2.37 78.22 $\pm$ 2.14 65.38
Co-MDA 64.56 $\pm$ 0.62 35.99 $\pm$ 1.21 79.66 $\pm$ 1.36 72.06 $\pm$ 1.66 66.33 $\pm$ 0.97 78.91 $\pm$ 1.87 66.34
WJDOT 89.06 $\pm$ 1.34 75.60 $\pm$ 1.84 89.99 $\pm$ 0.86 89.38 $\pm$ 0.77 85.32 $\pm$ 1.29 87.43 $\pm$ 1.23 86.13
WBT 92.38 $\pm$ 0.66 73.74 $\pm$ 1.07 88.89 $\pm$ 0.85 89.38 $\pm$ 1.26 85.53 $\pm$ 1.35 86.60 $\pm$ 1.63 86.09
DaDiL-R 91.97 $\pm$ 1.22 77.15 $\pm$ 1.32 85.41 $\pm$ 1.69 89.39 $\pm$ 1.03 84.49 $\pm$ 1.95 88.44 $\pm$ 1.29 86.14
DaDiL-E 90.45 $\pm$ 1.02 77.08 $\pm$ 1.21 86.79 $\pm$ 2.14 89.01 $\pm$ 1.35 84.04 $\pm$ 3.16 87.85 $\pm$ 1.06 85.87
GMM-WBT 92.23 $\pm$ 0.70 71.81 $\pm$ 1.78 84.72 $\pm$ 1.92 89.28 $\pm$ 1.55 87.51 $\pm$ 1.73 82.49 $\pm$ 1.81 84.67
GMM-DaDiL 91.72 $\pm$ 1.41 76.41 $\pm$ 1.89 89.68 $\pm$ 1.49 89.18 $\pm$ 1.17 86.05 $\pm$ 1.46 88.02 $\pm$ 1.12 86.85
(d) TEP.
Benchmarking Results. First, OT-based methods generally outperform other methods in MSDA. Overall, these methods solve a simpler task compared to deep DA methods, as they do not need to update the encoder network during adaptation. Second, the GMM-OT framework generally improves over using empirical OT. For instance, in the CWRU benchmark, GMM-WBT largely outperforms WBT [14, 13]. Furthermore, GMM-DaDiL outperforms its empirical counterpart on all benchmarks, as well as GMM-WBT. This point further showcases the power of dictionary learning in MSDA.
Figure 6: In (a), we analyse the performance of interpolations $\mathcal{B}((\lambda_{0},1-\lambda_{0});\mathcal{Q}_{S})$, $\mathcal{Q}_{S}=\{Q_{S_{1}},Q_{S_{2}}\}$ and $\mathcal{B}((\lambda_{0},1-\lambda_{0});\mathcal{P})$ with learned $\mathcal{P}=\{P_{1},P_{2}\}$ for GMM-WBT and GMM-DaDiL. In (b), we analyse the efficiency of barycenter-based methods under an increasing number of GMM components (number of samples for DaDiL and WBT). GMM-DaDiL has state-of-the-art performance even for the extreme case where $K=65$.
Better MSDA. We design an experiment to further showcase the contribution of dictionary learning for MSDA. We use the adaptation $(D,W)\rightarrow A$ in the Office-31 benchmark as the basis of our experiment. Note that GMM-DaDiL reconstructs the target domain via a barycenter $\mathcal{B}(\lambda_{T};\mathcal{P})$, where $\lambda_{T}$ and $\mathcal{P}$ are learned parameters. We thus ablate the learning of $\lambda_{T}$ and $\mathcal{P}$, i.e., we compare it to $\lambda_{T}=(\lambda_{0},1-\lambda_{0})$, $\lambda_{0}\in[0,1]$ and $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N_{S}}$. This generates a series of measures parametrized by $\lambda_{0}$. To further match $\mathcal{B}(\lambda_{T};\mathcal{Q}_{S})$ with $Q_{T}$, we transport it to $Q_{T}$ through eq. 15. Note that this corresponds to performing GMM-WBT with a barycenter calculated with $\lambda_{T}$. The overall experiment is shown in figure 6. While the performance of GMM-WBT remains approximately stable, that of GMM-DaDiL grows as we move closer to $\lambda_{T}^{\star}$ learned by dictionary learning (blue star). Overall, the interpolation space generated by atoms better captures the distributional shift occurring on the target domain.
Lighter MSDA. We further investigate the adaptation efficiency of our methods with respect the number of GMM components $n$. We use the adaptation task $(Cl,Pr,Rw)\rightarrow Ar$ from Office-Home to showcase our analysis. Note that this parameter is analogous to the number of samples in DaDiL or WBT. We thus show a comparison in figure 6 b. From this figure, we see that GMM-DaDiL surpasses all other methods over the entire range $n\in\{65,130,\cdots,910\}$, which showcases the advantage of using GMMs and dictionary learning. Especially, its empirical counterpart, DaDiL, needs a large number of samples in order to accurately represent probability measures.
Curiously, the performance of GMM-WBT and WBT are quite similar. Indeed, recent studies [18] show that Wasserstein barycenters are effective in compressing probability measures with respect the number of their samples. As a result, in this adaptation task, the GMM version of WBT has similar performance to the empirical version.
Visualization. In this experiment, we use the $(B,C)\rightarrow A$ adaptation task of CWRU. We are interested in visualizing the evolution of atoms and reconstructions with respect DaDiL and GMM-DaDiL iterations. We visualize this evolution through UMAP [11], i.e., we concatenate the data from DaDiL’s atoms, i.e., $\mathbf{x}_{i}^{(P_{c,it})}$, so that these are jointly embedded into $\mathbb{R}^{2}$. For GMM-DaDiL, we concatenate the mean parameters, i.e., $\mathbf{m}_{i}^{(P_{c,it})}$. We summarize our results in figure 7. Overall, as shown in figure 7 (a – d), GMM-DaDiL optimization is more stable than that of DaDiL, especially since we do not use mini-batches. This remark is also evidenced in the reconstructions in figure 7 (e – f).
(a) $P_{1,it}$
(b) $P_{2,it}$
(c) $\hat{P}_{1,it}$
(d) $\hat{P}_{2,it}$
(e) $\mathcal{B}(\lambda_{T};\mathcal{P}_{it})$
(f) $\mathcal{B}(\lambda_{T};\hat{\mathcal{P}}_{it})$
(g) $\mathcal{B}(\lambda_{T};\mathcal{P}^{\star})$
(h) $\mathcal{B}(\lambda_{T};\hat{\mathcal{P}}^{\star})$
Figure 7: From (a-d), we show the trajectory of atom distributions for GMM-DaDiL (a, b) and DaDiL (c, d). Blue and orange points represent the initializations and final vlaues for atoms at convergence. In (e, f), we show the trajectory of barycentric reconstructions for the target domain for these two methods. In (g, h), we show the reconstructions alongside target domain data at convergence.
5 Conclusion
In this work, we propose a novel framework for MSDA, using GMM-OT [5]. Especially, we propose a novel algorithm for calculating Wasserstein barycenters of GMMs (Algorithm 1). Based on this algorithm, we propose two new strategies for MSDA: GMM-WBT and GMM-DaDiL (Algorithm 2). The first method determines a labeled GMM on the target domain by transporting the barycenter of source domain GMMs towards the target. The second strategy uses dictionary learning to express each GMM in MSDA as the barycenter of learned GMMs. Overall, we propose methods that are lighter, better, faster than previous empirical OT methods in MSDA.

### v2

@@
 We empirically evaluate our proposed methods on four benchmarks in image classification and fault diagnosis, showing that we improve over the prior art while being faster and involving fewer parameters.
-Keywords:
-Domain Adaptation Optimal Transport Gaussian Mixture Models.
+Keywords Domain Adaptation  $\cdot$ Optimal Transport  $\cdot$ Gaussian Mixture Models.
 1 Introduction
@@
 Our contributions are threefold: 1.
-We propose a novel strategy for mapping the parameters of GMMs using OT (section 3.1 and theorem 3.1); 2.
+We propose a novel strategy for mapping the parameters of GMMs using OT (section 3.1 and theorem 1); 2.
 We propose a novel algorithm for computing mixture-Wasserstein barycenters of GMMs (algorithm 1 in section 3.3); 3.
 We propose an efficient parametric extension of the WBT and DaDiL algorithms based on GMMs (section 3.4).
-We highlight that, while GMMs were previously employed in single source DA [8], to the best of our knowledge this is the first work to leverage GMM-OT for MSDA.
+We highlight that, while GMMs were previously employed in single source DA [14, 15], to the best of our knowledge this is the first work to leverage GMM-OT for MSDA.
 Paper Organization.
@@
 As a result, one resorts to the minimization of the empirical risk, i.e., $\hat{h}=\text{argmin}_{h\in\mathcal{H}}\hat{\mathcal{R}}_{Q}(h)$, where ^$\hat{\mathcal{R}}_{Q}(h)=\dfrac{1}{n}\sum_{i=1}^{n}\mathcal{L}(h(\mathbf{x}_{i}^{(Q)}),y_{i}^{(Q)})$.
-From a theoretical standpoint, this framework is useful because $\mathcal{R}_{Q}$ is bounded by ^$\hat{\mathcal{R}}_{Q}$ and a complexity term depending on the number of samples $n$, and the Vapnik-Chervonenkis dimension of $\mathcal{H}$ [28, Section 6].
+From a theoretical standpoint, this framework is useful because $\mathcal{R}_{Q}$ is bounded by ^$\hat{\mathcal{R}}_{Q}$ and a complexity term depending on the number of samples $n$, and the Vapnik-Chervonenkis dimension of $\mathcal{H}$ [17, Section 6].
 As a result, $\hat{h}$ minimizing the empirical risk is guaranteed to generalize to unseen samples of $P$.
@@
 Optimal Transport is a field of mathematics concerned with the displacement of mass at least effort [18, 19].
-Given probability measures $P,Q\in\mathbb{P}(\mathcal{X})$, the Monge formulation [21, Section 2.2.] of OT seeks for a mapping $T$,
+Given probability measures $P,Q\in\mathbb{P}(\mathcal{X})$, the Monge formulation [18, Section 2.2.] of OT seeks for a mapping $T$,
 […]
@@
 Nonetheless, this problem poses technical difficulties, mainly due the constraint $T_{\sharp}P=Q$.
-A more tractable formulation was proposed by Kantorovich [21, Section 2.3.], and relies on OT plans,
+A more tractable formulation was proposed by Kantorovich [18, Section 2.3.], and relies on OT plans,
 […]
@@
 The following theorem provides us a strategy, [^1] We refer readers to our supplementary materials for all proofs and derivations.
-Theorem 3.1
+Theorem 1.
 Let $P$ and $Q$ be two GMMs with components $P_{i}=\mathcal{N}(\mathbf{m}_{i}^{(P)},(\mathbf{s}_{i}^{(P)})^{2})$ (resp.
@@
 where $\omega^{\star}$ is the solution of eq. 7.
-Equation 10 is similar to the barycentric mapping in Empirical Optimal Transport (EOT) [2, eq.
+Equation 10 is similar to the barycentric mapping in Empirical Optimal Transport (EOT) [6, eq.
 13], which serves as an approximation for the Monge mapping between $P$ and $Q$.
 In our case, the barycentric mappings act on the parameters of the GMM, rather than on its samples.
-Theorem 3.1 will be useful in the calculation of $\mathcal{MW}_{2}$ barycenters.
+Theorem 1 will be useful in the calculation of $\mathcal{MW}_{2}$ barycenters.
 3.2 Supervised Mixture-Wasserstein Distances
@@
 While simple, using an Euclidean distance for the soft-labels allows us to derive similar first-order conditions for $\mathcal{SMW}_{2}$,
-Theorem 3.2
-Under the same conditions of theorem 3.1, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{y}_{i}^{(P)}$ and $\mathbf{y}_{j}^{(Q)}$.
+Theorem 2.
+Under the same conditions of theorem 1, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{y}_{i}^{(P)}$ and $\mathbf{y}_{j}^{(Q)}$.
 The first order optimality conditions of $\mathcal{SMW}_{2}$ with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by eq.
@@
 As such, we adapt the definition of [21],
-Definition 1
+Definition 1.
 Given $C\geq 1$ GMMs $\mathcal{P}=\{P_{c}\}_{c=1}^{C}$ and a vector of barycentric coordinates $\lambda\in\Delta_{C}$, the $\mathcal{SMW}_{2}$ barycenter is given by,
@@
 When each $Q_{S_{\ell}}$ is a GMM, the parameters of $B$ are estimated through algorithm 1.
-Next, one solves for $\omega^{(T)}=\text{GMMOT}(B,Q_{T})$, so that the parameters of $B$ are transported towards $Q_{T}$ using theorems 3.1 and 3.2,
+Next, one solves for $\omega^{(T)}=\text{GMMOT}(B,Q_{T})$, so that the parameters of $B$ are transported towards $Q_{T}$ using theorems 1 and 2,
 $\hat{\mathbf{m}}_{i}^{(Q_{T})}=K_{B}\sum_{j=1}^{K_{T}}\omega_{ij}^{(T)}\mathbf{m}_{j}^{(Q_{T})}\text{, and, }\hat{\mathbf{s}}_{i}^{(Q_{T})}=K_{B}\sum_{j=1}^{K_{T}}\omega_{ij}^{(T)}\mathbf{s}_{j}^{(Q_{T})}.$ (15)
@@
 As a result, in this adaptation task, the GMM version of WBT has similar performance to the empirical version.
-Visualization.
-In this experiment, we use the $(B,C)\rightarrow A$ adaptation task of CWRU.
-We are interested in visualizing the evolution of atoms and reconstructions with respect DaDiL and GMM-DaDiL iterations.
-We visualize this evolution through UMAP [11], i.e., we concatenate the data from DaDiL’s atoms, i.e., $\mathbf{x}_{i}^{(P_{c,it})}$, so that these are jointly embedded into $\mathbb{R}^{2}$.
-For GMM-DaDiL, we concatenate the mean parameters, i.e., $\mathbf{m}_{i}^{(P_{c,it})}$.
-We summarize our results in figure 7.
-Overall, as shown in figure 7 (a – d), GMM-DaDiL optimization is more stable than that of DaDiL, especially since we do not use mini-batches.
-This remark is also evidenced in the reconstructions in figure 7 (e – f).
 (a) $P_{1,it}$
@@
 In (g, h), we show the reconstructions alongside target domain data at convergence.
+Visualization.
+In this experiment, we use the $(B,C)\rightarrow A$ adaptation task of CWRU.
+We are interested in visualizing the evolution of atoms and reconstructions with respect DaDiL and GMM-DaDiL iterations.
+We visualize this evolution through UMAP [32], i.e., we concatenate the data from DaDiL’s atoms, i.e., $\mathbf{x}_{i}^{(P_{c,it})}$, so that these are jointly embedded into $\mathbb{R}^{2}$.
+For GMM-DaDiL, we concatenate the mean parameters, i.e., $\mathbf{m}_{i}^{(P_{c,it})}$.
+We summarize our results in figure 7.
+Overall, as shown in figure 7 (a – d), GMM-DaDiL optimization is more stable than that of DaDiL, especially since we do not use mini-batches.
+This remark is also evidenced in the reconstructions in figure 7 (e – f).
 5 Conclusion
@@
 Overall, we propose methods that are lighter, better, faster than previous empirical OT methods in MSDA.
+Appendix A Introduction
+In this supplementary materials, we provide additional details about our theoretical and experimental results.
+Appendix B Proofs
+B.1 Proof for Theorem 1
+Theorem 3.
+Let $P$ and $Q$ be two GMMs with components $P_{i}=\mathcal{N}(\mathbf{m}_{i}^{(P)},(\mathbf{s}_{i}^{(P)})^{2})$ (resp.
+$Q_{j}$) and $\omega^{\star}$ be the solution of eq.
+10.
+The first-order optimality conditions of $\mathcal{MW}_{2}^{2}$, with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by,
+$\mathbf{m}_{i}=T_{\omega}(\mathbf{m}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{m}_{j}^{(Q)}\text{, and }\mathbf{s}_{i}=T_{\omega}(\mathbf{s}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{s}_{j}^{(Q)}.$ (17)
+Proof.
+Under the GMM-OT framework, the $2-$mixture Wasserstein distance is given by,
+$\mathcal{MW}_{2}(P,Q)^{2}=\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathcal{W}_{2}(P_{i},Q_{j})^{2},$
+where $\omega_{ij}^{\star}$ is the mass transported from component $i$ in $P$ to component $j$ in $Q$ by the solution to the GMM-OT problem.
+Since $P_{i}$ and $Q_{j}$ are axis-aligned Gaussian measures,
+$\mathcal{W}_{2}(P_{i},Q_{j})^{2}=\lVert\mathbf{m}_{i}-\mathbf{m}_{j}^{(Q)}\rVert_{2}^{2}+\lVert\mathbf{s}_{i}-\mathbf{s}_{j}^{(Q)}\rVert_{2}^{2},$
+where $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are the free parameters we want to determine.
+Henceforth we derive the first-order conditions for $\mathbf{m}_{i}$.
+The derivations for $\mathbf{s}_{i}$ are completely analogous.
+By linearity,
+$\dfrac{\partial\mathcal{MW}_{2}^{2}}{\partial\mathbf{m}_{i}}=2\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}(\mathbf{m}_{i}-\mathbf{m}_{j}^{(Q)})=2\biggr{(}p_{i}\mathbf{m}_{i}-\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathbf{m}_{j}^{(Q)}\biggr{)},$
+equating this last term to 00, one gets the desired equality. ∎
+The result of theorem 3 is quite similar to the barycentric mapping of [33, 6] which was previously used to map points in $P$ to points in $Q$. Here, we transport the parameters of GMMs instead.
+B.2 Proof of Theorem 2
+Theorem 4.
+Under the same conditions of theorem 3, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{y}_{i}^{(P)}$ and $\mathbf{y}_{j}^{(Q)}$.
+The first order optimality conditions of $\mathcal{SMW}_{2}$ with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by eq.
+17.
+Furthermore, for $\mathbf{y}_{i}$,
+$\mathbf{y}_{i}=T_{\omega}(\mathbf{y}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{y}_{j}^{(Q)}.$ (18)
+Proof.
+Note that the label distance term added to $\mathcal{SMW}$ does not interfere with the terms on $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$, hence the optimality conditions with respect these variables remain unchanged.
+Due the independence of these terms, the first-order optimality condition with respect $\mathbf{y}_{i}$ is,
+$\dfrac{\partial\mathcal{SMW}_{2}^{2}}{\partial\mathbf{y}_{i}}=2\beta\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}(\mathbf{y}_{i}-\mathbf{y}_{j}^{(Q)})=2\beta\biggr{(}p_{i}\mathbf{y}_{i}-\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathbf{y}_{j}^{(Q)}\biggr{)},$
+which, for $\beta>0$, is zero if and only if $\mathbf{y}_{i}=T_{\omega}(\mathbf{y}_{i}^{(P)})$. ∎
+B.3 Derivation of Algorithm 1
+In this section, we derive the first-order conditions for the functional $B\mapsto\sum_{c}\lambda_{c}\mathcal{SMW}_{2}(B,P_{c})^{2}$.
+Let $B$ be parametrized by $\theta_{B}=\{(\mathbf{m}_{i}^{(B)},\mathbf{s}_{i}^{(B)},\mathbf{y}_{i}^{(B)})\}_{i=1}^{K_{B}}$, where these variables are initialized randomly as in algorithm 1.
+For fixed $\theta_{B}$, we find $\omega_{1}^{\star},\cdots,\omega_{C}^{\star}$ transport plans.
+Then, for fixed transport plans, we solve,
+$\underset{\theta_{B}}{\text{argmin}}\,\mathcal{L}(\theta_{B})$ $=\sum_{c=1}^{C}\lambda_{c}\sum_{i=1}^{K_{B}}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{(\star)}M_{c,i,j},$
+where $\text{where }M_{c,i,j}$ $=\lVert\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})}\rVert_{2}^{2}+\lVert\mathbf{s}_{i}^{(B)}-\mathbf{s}_{j}^{(P_{c})}\rVert_{2}^{2}+\beta\lVert\mathbf{y}_{i}^{(B)}-\mathbf{y}_{j}^{(P_{c})}\rVert_{2}^{2}$
+which can be optimized by taking derivatives with respect $\mathbf{m}^{(B)}_{i}$, $\mathbf{s}^{(B)}_{i}$ and $\mathbf{y}^{(B)}_{i}$.
+For instance, taking the derivative of $\mathcal{L}(\theta_{B})$ with respect $\mathbf{m}_{i}^{(B)}$,
+$\dfrac{\partial\mathcal{L}}{\partial\mathbf{m}_{i}^{(B)}}=2\sum_{c=1}^{C}\lambda_{k}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{\star}(\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})})=\dfrac{2}{K_{B}}\mathbf{m}_{i}^{(B)}-2\sum_{c=1}^{C}\lambda_{c}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{\star}\mathbf{m}_{j}^{(P_{c})}$
+setting the derivative to 00, one has,
+$\mathbf{m}_{i}^{(B)}$ $=\sum_{c=1}^{C}\lambda_{c}T_{\omega_{c}^{\star}}(\mathbf{m}_{i}^{(B)}).$
+Similar results can be acquired for $\mathbf{s}_{i}^{(B)}$ and $\mathbf{y}_{i}^{(B)}$ by taking the appropriate derivatives.
+Appendix C Additional Experiments
+Figure 8: Evolution of barycentric coordinates over the simplex $\Delta_{C}$.
+Toy Example.
+We analyze the evolution of the atoms in GMM-DaDiL as training progresses (figure 9).
+Note that, initially, the means are close to each other, so there are no distinction between classes.
+Likewise, in our initialization, the labels are uniform histograms over the classes, so that the atoms are not certain about which class they represent.
+With 50 iterations, DaDiL manages to spread the class clusters, but, at this point, the barycentric coordinates of each domain are still too similar (i.e., $\lambda_{\ell}\approx\mathds{1}_{K}/K$, see figure 2.a in the main paper).
+As the training progresses, the barycentric coordinates of the different domains begins to spread over the simplex $\Delta_{C}$ (see figure 8).
+As a result, the atoms begin to be different from each other (compare figures 9 (b) and (c)).
+Hyper-parameter Setting: Here we detail the hyper-parameter setting for our method.
+GMM-WBT involves a single hyper-parameter: the number of components $K$ in the GMMs.
+DaDiL involves a few other parameters: the number of atoms $C$, the learning rate $\eta$ and the number of optimization steps $N_{iter}$.
+As [12], we use Adam [34] for optimizing eq.
+16 (see main paper).
+Over all our experiments we set $\eta=10^{-1}$.
+We show our hyper-parameters in table 3.
+We remark that, as we analyze in Fig 6.
+b (main paper), our proposed methods, especially GMM-DaDiL, are robust to the choice of hyper-parameters.
+Table 3: Hyper-parameter setting
+Benchmark $K$ (GMM-WBT) $K$ (GMM-DaDiL) $C$ $N_{iter}$
+Office 31 155 186 3 150
+Office-Home 910 455 4 150
+TEP 261 174 7 150
+CWRU 30 90 5 200
+(a) 1st iteration.
+(b) 50th iteration.
+(c) 100th iteration.
+(d) 150th iteration.
+(e) 200th iteration.
+Figure 9: Evolution of atoms as the training progresses.

### v3

@@
 Abstract
-In this paper, we tackle Multi-Source Domain Adaptation (MSDA), a task in transfer learning where one adapts multiple heterogeneous, labeled source probability measures towards a different, unlabeled target measure.
-We propose a novel framework for MSDA, based on Optimal Transport (OT) and Gaussian Mixture Models (GMMs).
-Our framework has two key advantages.
-First, OT between GMMs can be solved efficiently via linear programming.
-Second, it provides a convenient model for supervised learning, especially classification, as components in the GMM can be associated with existing classes.
-Based on the GMM-OT problem, we propose a novel technique for calculating barycenters of GMMs.
-Based on this novel algorithm, we propose two new strategies for MSDA: GMM-WBT and GMM-DaDiL.
-We empirically evaluate our proposed methods on four benchmarks in image classification and fault diagnosis, showing that we improve over the prior art while being faster and involving fewer parameters.
+[…]
 Keywords Domain Adaptation  $\cdot$ Optimal Transport  $\cdot$ Gaussian Mixture Models.
@@
 For MSDA, OT has been used for aggregating the multiple source domains into a barycentric domain [8, 9], which is later transported to the target domain, or by weighting source domain measures [10].
-This work considers the problem of Wasserstein Dictionary Learning (WDL), which was initially proposed by [11] for histogram data.
+Our work considers the problem of Wasserstein Dictionary Learning (WDL), initially proposed by [11] for histogram data.
 This problem was later generalized by [12], for empirical measures, which allowed its application to MSDA.
-In [12], one learns to express domains in MSDA as a barycenter of atom measures, which have a free, learnable support.
+In [12], one expresses domains in MSDA as a barycenter of atom measures, which have a free, learnable support.
 As a result, the work of [12] learns how to interpolate distributional shift between the measures in MSDA.
-However, previous algorithms relying on Wasserstein barycenters, such as Wasserstein Barycenter Transport (WBT) [8, 9] and Dataset Dictionary Learning (DaDiL) [12], are limited in scale, since the number of points the support of the empirical measures scale with the number of samples in the original datasets.
+[…]
+(a) GMM-WBT
+[…]
+(b) GMM-DaDiL
+[…]
+However, previous algorithms relying on Wasserstein barycenters, such as WBT [8, 9] and DaDiL [12], are limited in scale, since the number of points the support of the empirical measures scale with the number of samples in the original datasets.
 As a consequence, previous works such as [8, 9] are limited to small scale datasets, or rely on mini-batch optimization [12], which introduces artifacts in the OT.
-To tackle the aforementioned limitations, in this paper we propose a novel, parametric framework for barycentric-based MSDA algorithms.
-based on recent developments in OT between Gaussian Mixture Models (GMMs) [13].
+To tackle these limitations, in this paper we propose a novel, parametric framework for barycentric-based MSDA algorithms.
+based on OT between GMMs [13].
 We present an overview of our methods in figure 1.
-(a) GMM-WBT
-(b) GMM-DaDiL
-Figure 1: Overview of proposed methods.
-Squares represent actual domains, circles represent barycenters and triangles represent learned measures.
-Blue and orange elements represent labeled and unlabeled measures respectively.
-In GMM-WBT, a labeled GMM is determined for the target domain by transporting the barycenter of sources.
-In GMM-DaDiL, we learn to express each domain as a barycenter of learned GMMs, called atoms, through dictionary learning.
-Summary of Contributions.
 Our contributions are threefold: 1.
@@
 We highlight that, while GMMs were previously employed in single source DA [14, 15], to the best of our knowledge this is the first work to leverage GMM-OT for MSDA.
-Paper Organization.
 The rest of this paper is divided as follows.
@@
 2 Preliminaries
-In this section, we cover the necessary background for our paper. We divide our discussion into Gaussian mixtures, domain adaptation, optimal transport and dataset dictionary learning.
-Gaussian Mixtures.
-We denote the set of probability measures over a set $\mathcal{X}$ as $\mathbb{P}(\mathcal{X})$.
-A Gaussian measure corresponds to $P_{\theta}\in\mathbb{P}(\mathcal{X})$ with the following density,
+2.1 Gaussian Mixtures
+We denote the set of probability measures over a set $\mathcal{X}$ as $\mathbb{P}(\mathcal{X})$. A Gaussian measure corresponds to $P_{\theta}\in\mathbb{P}(\mathcal{X})$ with density,
 $f_{\theta}(\mathbf{x})$ $=\dfrac{1}{\sqrt{(2\pi)^{d}\det(\mathbf{C}^{(P)})}}\exp\biggr{(}-\dfrac{1}{2}(\mathbf{x}-\mathbf{m}^{(P)})^{T}(\mathbf{C}^{(P)})^{-1}(\mathbf{x}-\mathbf{m}^{(P)})\biggr{)},$
@@
 $P_{\theta}=\sum_{k=1}^{K}p_{k}P_{k}\text{, where }P_{k}=\mathcal{N}(\mathbf{m}_{k}^{(P)},\mathbf{C}_{k}^{(P)})\text{, and }\mathbf{p}\in\Delta_{K},$ (1)
-where $\Delta_{K}=\{\mathbf{p}\in\mathbb{R}^{K}:\sum_{k=1}^{K}p_{k}=1\}$.
+where $\Delta_{K}=\{\mathbf{p}\in\mathbb{R}^{K}_{+}:\sum_{k=1}^{K}p_{k}=1\}$.
 Following [13], we denote the subset of $\mathbb{P}(\mathbb{R}^{d})$ of probability measures which can be written as Gaussian mixtures with less than $K$ components by $\text{GMM}_{d}(K)$, and $\text{GMM}_{d}(\infty)=\cup_{k\geq 0}\text{GMM}_{d}(K)$.
@@
 $\theta^{\star}$ $=\underset{\theta\in\Theta}{\text{argmax}}\,\sum_{i=1}^{n}\log P_{\theta}(\mathbf{x}_{i}^{(P)}),$ (2)
-where $\Theta=\{\{p_{k},\mathbf{m}_{k}^{(P)},\mathbf{C}_{k}^{(P)}\}_{k=1}^{K}:\sum_{k}p_{k}=1\text{ and }\mathbf{C}_{k}^{(P)}\in\mathbb{S}^{d}\}$.
+where $\Theta=\{\{p_{k},\mathbf{m}_{k}^{(P)},\mathbf{C}_{k}^{(P)}\}_{k=1}^{K}:\mathbf{m}_{k}^{(P)}\in\mathbb{R}^{d}\text{ and }\mathbf{C}_{k}^{(P)}\in\mathbb{S}^{d}\}$.
 While equation 2 has no closed-form solution, one can solve this optimization problem through the celebrated Expectation-Maximization (EM) algorithm [16].
-Learning Theory and Domain Adaptation.
+2.2 Domain Adaptation
 In this paper, we focus on the problem of classification.
-Given a feature space $\mathcal{X}=\mathbb{R}^{d}$ and a label space $\mathcal{Y}=\{1,\cdots,n_{cl}\}$, this problem corresponds to finding $h\in\mathcal{H}\subset\mathcal{X}^{\mathcal{Y}}$ that correctly classifies data $\{(\mathbf{x}_{i}^{(Q)},y_{i}^{(Q)})\}_{i=1}^{n}$.
-In this paper, we use the Empirical Risk Minimization (ERM) framework [17], as it is useful for domain adaptation theory.
-As follows, one assumes $\mathbf{x}_{i}^{(Q)}\overset{iid}{\sim}Q$, for a measure $Q\in\mathbb{P}(\mathcal{X})$, and $h_{0}\in\mathcal{X}^{\mathcal{Y}}$ such that $y_{i}^{(Q)}=h_{0}(\mathbf{x}_{i}^{(Q)})$.
+Given a feature space $\mathcal{X}=\mathbb{R}^{d}$ and a label space $\mathcal{Y}=\{1,\cdots,n_{cl}\}$, this problem corresponds to finding $h\in\mathcal{H}\subset\mathcal{Y}^{\mathcal{X}}$ that correctly classifies data $\{(\mathbf{x}_{i}^{(Q)},y_{i}^{(Q)})\}_{i=1}^{n}$.
+We use the Empirical Risk Minimization (ERM) framework [17], as it is useful for domain adaptation theory.
+As follows, one assumes $\mathbf{x}_{i}^{(Q)}\overset{iid}{\sim}Q$, for a measure $Q\in\mathbb{P}(\mathcal{X})$, and $h_{0}:\mathcal{X}\rightarrow\mathcal{Y}$ such that $y_{i}^{(Q)}=h_{0}(\mathbf{x}_{i}^{(Q)})$.
 $h_{0}$ is called ground-truth labeling function.
@@
 From a theoretical standpoint, this framework is useful because $\mathcal{R}_{Q}$ is bounded by ^$\hat{\mathcal{R}}_{Q}$ and a complexity term depending on the number of samples $n$, and the Vapnik-Chervonenkis dimension of $\mathcal{H}$ [17, Section 6].
-As a result, $\hat{h}$ minimizing the empirical risk is guaranteed to generalize to unseen samples of $P$.
-Nevertheless, the assumption that unseen examples come from a fixed measure $P$ is seldom verified in practice [1], since the conditions upon which data is acquired may change.
+As a result, $\hat{h}$ minimizing the empirical risk is guaranteed to generalize to unseen samples of $Q$.
+Nevertheless, the assumption that unseen examples come from a fixed measure $Q$ is seldom verified in practice [1], since the conditions upon which data is acquired may change.
 In this case, models are required to adapt to new data, but at the same time re-training a model from the scratch is likely costly and data intensive.
@@
 Our goal is to learn a classifier on $Q_{T}$ by leveraging the knowledge from the source domains.
+2.3 Optimal Transport
 Optimal transport is a field of mathematics concerned with the displacement of mass at least effort [18, 19].
@@
 […]
-where $T_{\sharp}$ is the push-forward mapping of $T$, i.e., $T_{\sharp}P(A)=P(T^{-1}(A))$.
+where $T_{\sharp}$ is the push-forward mapping of $T$, i.e., $T_{\sharp}P(A)=P(T^{-1}(A))$, and $c:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a ground-cost, that is, a measure of transportation effort.
 Nonetheless, this problem poses technical difficulties, mainly due the constraint $T_{\sharp}P=Q$.
@@
 Under these conditions, for $\mathbf{C}^{(P)}=\mathbf{S}^{(P)}(\mathbf{S}^{(P)})^{T}$,
-$\mathcal{W}_{2}^{2}(P,Q)=\lVert\mathbf{m}^{(P)}-\mathbf{m}^{(Q)}\rVert^{2}_{2}+\text{Tr}\biggr{(}\mathbf{C}^{(P)}+\mathbf{C}^{(Q)}-2(\mathbf{S}^{(P)}\mathbf{C}^{(Q)}\mathbf{S}^{(P)})^{\nicefrac{{1}}{{2}}}\biggr{)},$
+$\mathcal{W}_{2}(P,Q)^{2}=\lVert\mathbf{m}^{(P)}-\mathbf{m}^{(Q)}\rVert^{2}_{2}+\text{Tr}\biggr{(}\mathbf{C}^{(P)}+\mathbf{C}^{(Q)}-2(\mathbf{S}^{(P)}\mathbf{C}^{(Q)}\mathbf{S}^{(P)})^{\nicefrac{{1}}{{2}}}\biggr{)},$
 This expression can be further simplified for axis-aligned Gaussians, i.e., $\mathbf{S}^{(P)}=\text{diag}(\mathbf{s}^{(P)})$, with $\mathbf{s}^{(P)}\in\mathbb{R}^{d}_{+}$,
-$\mathcal{W}_{2}^{2}(P,Q)$ $=\lVert\mathbf{m}^{(P)}-\mathbf{m}^{(Q)}\rVert^{2}_{2}+\lVert\mathbf{s}^{(P)}-\mathbf{s}^{(Q)}\rVert_{2}^{2}.$ (6)
+$\mathcal{W}_{2}(P,Q)^{2}$ $=\lVert\mathbf{m}^{(P)}-\mathbf{m}^{(Q)}\rVert^{2}_{2}+\lVert\mathbf{s}^{(P)}-\mathbf{s}^{(Q)}\rVert_{2}^{2}.$ (6)
 Henceforth, we assume axis-aligned Gaussian measures.
-In this paper, we use the GMM-OT framework of [13], which is convenient to our setting for 2 reasons.
+Remark.
+Here, we give further insight into the hypothesis of using axis-aligned Gaussian measures.
+We use this assumption for numerical stability purposes, i.e., estimating the covariance matrix of GMMs in high dimensions is much more difficult than estimating the standard deviation vector $\mathbf{s}^{(P)}$.
+Here, one has two choices.
+First, it is possible to introduce a transformation so as to force features to be uncorrelated (e.g., through principal components analysis).
+This approach, nonetheless, requires more data points per domain than features, which is not always feasible.
+Conversely, one can increase the number of components for expressing the shape of the data (see Fig.
+2).
+As we show in our experiments sections, we achieve good adaptation performance, even while sampling points from axis-aligned GMMs.
+Figure 2: Illustration of axis-aligned GMMs. This hypothesis leads to GMMs that need more components to express the underlying data distribution.
+We use the GMM-OT framework of [13], which is convenient to our setting for 2 reasons.
 First, they are able to represent measures with sub-populations, such as those commonly encountered in classification and domain adaptation.
@@
 $\mathcal{MW}_{2}(P,Q)^{2}$ $=\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathcal{W}_{2}(P_{i},Q_{j})^{2}.$ (8)
-Dataset Dictionary Learning [12] is an OT-based framework for expressing probability measures as barycenters of synthetic measures, called atoms.
+DaDiL [12] is an OT-based framework for expressing probability measures as barycenters of synthetic measures, called atoms.
+In this case, the authors use empirical measures, i.e., $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{(\mathbf{x}_{i}^{(P)},\mathbf{y}_{i}^{(P)})}$.
 The framework is inspired by dictionary learning literature [11].
-The authors introduce atoms $\mathcal{P}=\{\hat{P}_{c}\}_{c=1}^{C}$ and barycentric coordinates $\Lambda=\{\lambda_{\ell}\}_{\ell=1}^{N}$, such that each measure in MSDA is expressed as a Wasserstein barycenter $\hat{Q}_{\ell}=\mathcal{B}(\lambda_{\ell};\mathcal{P})$.
+The authors introduce atoms $\mathcal{P}=\{\hat{P}_{c}\}_{c=1}^{C}$ and barycentric coordinates $\Lambda=\{\lambda_{\ell}\}_{\ell=1}^{N}$, such that each measure in MSDA is expressed as a Wasserstein barycenter $\hat{Q}_{\ell}=\mathcal{B}(\lambda_{\ell},\mathcal{P})$.
 This framework leads to the following optimization problem,
-$\mathcal{L}(\Lambda,\mathcal{P})$ $=\underset{\Lambda,\mathcal{P}}{\text{argmin}}\,\mathcal{W}_{2}(\hat{Q}_{T},\mathcal{B}(\lambda_{\ell};\mathcal{P}))+\sum_{\ell=1}^{N}\mathcal{W}_{c,2}(\hat{Q}_{\ell},\mathcal{B}(\lambda_{\ell};\mathcal{P}))^{2},$ (9)
+$(\Lambda^{\star},\mathcal{P}^{\star})$ $=\underset{\Lambda,\mathcal{P}}{\text{argmin}}\,\mathcal{W}_{2}(\hat{Q}_{T},\mathcal{B}(\lambda_{\ell};\mathcal{P}))+\sum_{\ell=1}^{N}\mathcal{W}_{c,2}(\hat{Q}_{\ell},\mathcal{B}(\lambda_{\ell};\mathcal{P}))^{2},$ (9)
 where $c\biggr{(}(\mathbf{x}^{(Q_{\ell})},\mathbf{y}^{(Q_{\ell})}),(\mathbf{x}^{(B_{\ell})},\mathbf{y}^{(B_{\ell})})\biggr{)}=\lVert\mathbf{x}^{(Q_{\ell})}-\mathbf{x}^{(B_{\ell})}\rVert_{2}^{2}+\beta\lVert\mathbf{y}^{(Q_{\ell})}-\mathbf{y}^{(B_{\ell})}\rVert_{2}^{2}$, and $\beta\geq 0$ is a constant expressing how costly it is to move samples from different classes.
 This framework makes it easier to express the distributional shift between the different $\mathcal{Q}=\{\hat{Q}_{1},\cdots,\hat{Q}_{N_{S}},\hat{Q}_{T}\}$.
-Especially, since $\hat{B}_{\ell}=\mathcal{B}(\lambda_{\ell};\mathcal{P})$ is labeled, one can synthesize labeled target domain data by reconstructing the target measure with $\lambda_{T}:=\lambda_{N+1}$.
+Especially, since $\hat{B}_{\ell}=\mathcal{B}(\lambda_{\ell},\mathcal{P})$ is labeled, one can synthesize labeled target domain data by reconstructing the target measure with $\lambda_{T}:=\lambda_{N+1}$.
 3 Methodological Contributions
 3.1 First Order Analysis of $\mathcal{MW}_{2}$
-In this section we analyze $P\mapsto\mathcal{MW}_{2}(P,Q)$, for a fixed $Q$.
+In this section we analyze $P\mapsto\mathcal{MW}_{2}(P,Q)^{2}$, for a fixed $Q$.
 We are particularly interested on how to map the components of $P$ towards $Q$, while minimizing this distance.
-The following theorem provides us a strategy, [^1] We refer readers to our supplementary materials for all proofs and derivations.
+The following theorem provides us a strategy,
 Theorem 1.
@@
 $Q_{j}$) and $\omega^{\star}$ be the solution of eq.
-8.
+7.
 The first-order optimality conditions of $\mathcal{MW}_{2}^{2}$, with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by,
-$\mathbf{m}_{i}=T_{\omega^{\star}}(\mathbf{m}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{m}_{j}^{(Q)}\text{, and }\mathbf{s}_{i}=T_{\omega^{\star}}(\mathbf{s}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{s}_{j}^{(Q)},$ (10)
+$\hat{\mathbf{m}}_{i}=T_{\omega^{\star}}(\mathbf{m}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{m}_{j}^{(Q)}\text{, and }\hat{\mathbf{s}}_{i}=T_{\omega^{\star}}(\mathbf{s}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{s}_{j}^{(Q)},$ (10)
 where $\omega^{\star}$ is the solution of eq. 7.
+Proof.
+Our proof relies on the analysis of $\{(\mathbf{m}_{i},\mathbf{s}_{i})\}_{i=1}^{K_{P}}\mapsto\mathcal{MW}_{2}(P,Q)^{2}$ (c.f., equation 8). Given $\omega^{\star}=\text{GMMOT}(P,Q)$,
+$\dfrac{\partial\mathcal{MW}_{2}^{2}}{\partial\mathbf{m}_{i}}=2\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}(\mathbf{m}_{i}-\mathbf{m}_{j}^{(Q)})=2\biggr{(}p_{i}\mathbf{m}_{i}-\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathbf{m}_{j}^{(Q)}\biggr{)},$
+and, by equating this last term to 00, one gets the desired equality. ∎
 Equation 10 is similar to the barycentric mapping in Empirical Optimal Transport (EOT) [6, eq.
@@
 As such, it is necessary to equip the components of GMMs with labels that represent the classes in the datasets.
-We propose doing so through a simple heuristic, based on the law of total probability,
-$P(y|k)$ $=\sum_{i=1}^{n}P_{\theta}(\mathbf{x}_{i}|k)P(y|\mathbf{x}_{i})=\sum_{i=1}^{n}\biggr{(}\dfrac{p_{k}P_{k}(\mathbf{x}_{i})}{\sum_{k^{\prime}}p_{k^{\prime}}P_{k^{\prime}}(\mathbf{x}_{i})}\biggr{)}P(y|\mathbf{x}_{i}),$ (11)
-where $P_{\theta}(\mathbf{x}_{i}|k)$ corresponds to the probability that $\mathbf{x}_{i}$ comes form the $k-$th component of the GMM.
-Henceforth, we denote the soft-labels of components by $\mathbf{y}_{i}^{(P)}$.
-We may use the labeled GMM for obtaining a classifier, by sampling $\mathbf{x}\sim P_{\theta}$.
-If $\mathbf{x}$ is sampled from the $k-$th component, we simply assign $P(y|k)$ as its soft-label.
-We illustrate our strategy in figure 2.
-(a) Gaussian Mixture $P_{\theta}$
-(b) $\text{argmax}_{k}P_{\theta}(\mathbf{x}|k)$.
-(c) Labeled GMM
-Figure 2: Illustration of the labeling process of Gaussian mixtures.
-(a) First, the mixture $P_{\theta}$ is learned.
-(b) Then, one estimates $P_{\theta}(\mathbf{x}|k)$ for each $\mathbf{x}_{i}$.
-(c) The label of each component can be estimated using the law of total probability.
+We propose doing so through a simple heuristic, especially, we model $P(\mathbf{x}|y)$ through a GMM.
+We then concatenate the $n_{c}$ obtained GMMs, and assign, for the $k-$th GMM of the $y-$th class, $v_{k,y^{\prime}}^{(P)}=\delta(y^{\prime}-y)$, i.e., a vector of $n_{c}$ components, and $1$ on the $y-$th entry.
+We can assure that the resulting weights sum to $1$ by dividing their value by $\sum_{y=1}^{n_{c}}\sum_{k=1}^{K}p_{k,y}$, where $p_{k,y}$ corresponds to the weight of the $k-$th component of the $y-$th GMM.
+Given a GMM $\{p_{k},\mathbf{m}_{k}^{(P)},\mathbf{s}_{k}^{(P)},\mathbf{v}_{k}^{(P)}\}_{k=1}^{K}$, we define a classifier through Maximum a Posteriori (MAP) estimation.
+This strategy is carried out through,
+$\hat{h}_{MAP}(\mathbf{x})=\underset{y=1,\cdots,n_{c}}{\text{argmax}}\,P(y|\mathbf{x})=\sum_{k=1}^{K}\underbrace{P_{\theta}(k|\mathbf{x})}_{p_{k}P_{k}(\mathbf{x})/\sum_{k^{\prime}}p_{k^{\prime}}P_{k^{\prime}}(\mathbf{x})}\underbrace{P(y|k)}_{v_{k,y}^{(P)}},$ (11)
+we use this classifier in a few illustrative examples in section 4.3.
+Remark.
+In equation 11, we are implicitly assuming that the component $k$ is conditionally independent with $y$ given $\mathbf{x}$.
+This remark is intuitive, as $\mathbf{x}$ explains, at the same time, the component and the label.
 Similarly to EOT, when the mixtures $P$ and $Q$ are labeled, one needs to take into account the labels in the ground-cost. Given $\beta>0$, we propose the following distance between labeled GMMs,
-$\mathcal{SMW}_{2}(P,Q)^{2}$ $=\underset{\omega\in\Gamma(\mathbf{p},\mathbf{q})}{\text{argmin}}\,\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}(\mathcal{W}_{2}(P_{i},Q_{j})^{2}+\beta\lVert\mathbf{y}_{i}^{(P)}-\mathbf{y}_{j}^{(Q)}\rVert_{2}^{2}).$ (12)
+$\mathcal{SMW}_{2}(P,Q)^{2}$ $=\underset{\omega\in\Gamma(\mathbf{p},\mathbf{q})}{\text{min}}\,\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}(\mathcal{W}_{2}(P_{i},Q_{j})^{2}+\beta\lVert\mathbf{v}_{i}^{(P)}-\mathbf{v}_{j}^{(Q)}\rVert_{2}^{2}).$ (12)
 While simple, using an Euclidean distance for the soft-labels allows us to derive similar first-order conditions for $\mathcal{SMW}_{2}$,
 Theorem 2.
-Under the same conditions of theorem 1, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{y}_{i}^{(P)}$ and $\mathbf{y}_{j}^{(Q)}$.
+Under the same conditions of theorem 1, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{v}_{i}^{(P)}$ and $\mathbf{v}_{j}^{(Q)}$.
 The first order optimality conditions of $\mathcal{SMW}_{2}$ with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by eq.
 10.
-Furthermore, for $\mathbf{y}_{i}$,
-$\hat{\mathbf{y}}_{i}=T_{\omega}(\mathbf{y}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{y}_{j}^{(Q)}.$ (13)
+Furthermore, for $\mathbf{v}_{i}$,
+$\hat{\mathbf{v}}_{i}=T_{\omega}(\mathbf{v}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{v}_{j}^{(Q)}.$ (13)
+Proof.
+The label distance term in $\mathcal{SMW}$ is independent of $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$, hence the optimality conditions of these variables remain unchanged.
+Therefore, the first-order optimality condition with respect $\mathbf{v}_{i}$ is,
+$\dfrac{\partial\mathcal{SMW}_{2}^{2}}{\partial\mathbf{v}_{i}}=2\beta\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}(\mathbf{v}_{i}-\mathbf{v}_{j}^{(Q)})=2\beta\biggr{(}p_{i}\mathbf{v}_{i}-\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathbf{v}_{j}^{(Q)}\biggr{)},$
+which, for $\beta>0$, is zero if and only if $\mathbf{v}_{i}=T_{\omega}(\mathbf{v}_{i}^{(P)})$. ∎
+Remark.
+In equation 12, we are heuristically adding a label regularization term to the $\mathcal{MW}_{2}$ distance.
+The actual continuous counterpart (between samples, rather than components) is currently beyond the scope of this paper, but methodologically, this choice remains valid, and is closer to the contributions of [chen2018optimal].
 3.3 Mixture Wasserstein Barycenters
-In this section, we detail our first methodological contribution: a new algorithm for computing barycenters of GMMs under the $\mathcal{MW}_{2}$ and $\mathcal{SMW}_{2}$ metrics.
-As such, we adapt the definition of [21],
+In this section, we detail a new algorithm for computing barycenters of GMMs under the $\mathcal{MW}_{2}$ and $\mathcal{SMW}_{2}$ metrics. As such, we adapt the definition of [21],
 Definition 1.
-Given $C\geq 1$ GMMs $\mathcal{P}=\{P_{c}\}_{c=1}^{C}$ and a vector of barycentric coordinates $\lambda\in\Delta_{C}$, the $\mathcal{SMW}_{2}$ barycenter is given by,
-$P^{\star}=\mathcal{B}(\lambda,\mathcal{P})=\underset{B\in\text{GMM}_{d}(\infty)}{\text{argmin}}\,\biggr{\{}\mathcal{L}(B)=\sum_{c=1}^{C}\lambda_{c}\mathcal{SMW}_{2}(B,P_{c})^{2}\biggr{\}}.$ (14)
+Given $C\geq 1$ GMMs $\mathcal{P}=\{P_{c}\}_{c=1}^{C}$, $K_{B}\geq 1$, and a vector of barycentric coordinates $\lambda\in\Delta_{C}$, the $\mathcal{SMW}_{2}$ barycenter is given by,
+$B^{\star}=\mathcal{B}(\lambda,\mathcal{P})=\underset{B\in\text{GMM}_{d}(K_{B})}{\text{argmin}}\,\biggr{\{}\mathcal{L}(B)=\sum_{c=1}^{C}\lambda_{c}\mathcal{SMW}_{2}(B,P_{c})^{2}\biggr{\}}.$ (14)
 When the GMMs in $\mathcal{P}$ are unlabeled, one may define, by analogy, a barycenter under the $\mathcal{MW}_{2}$.
@@
 Inspired by previous results in empirical Wasserstein barycenters [22, 12], we propose a novel strategy for computing $\mathcal{B}(\lambda,\mathcal{P})$.
-Our idea consists on initializing the parameters $\{(\mathbf{m}_{i}^{(B)},\mathbf{s}_{i}^{(B)},\mathbf{y}_{i}^{(B)})\}_{i=1}^{K_{B}}$ at random.
-We then calculate $C$ GMM-OT transport plans $\{\omega^{(c)}\}_{c=1}^{C}$ between $P_{c}$ and $B$.
-Next, we use equations 10 and 13 to update the parameters of the barycenter $B$.
-We refer readers to our supplementary materials for further details.
-This strategy effectively minimizes $\mathcal{L}$ in eq.
-14.
-We show our strategy in Algorithm 1.
-1 function mixture_wasserstein_barycenter($\{\mathbf{M}^{(P_{c})}\}_{c=1}^{C}$, $\{\mathbf{S}^{(P_{c})}\}_{c=1}^{C}$, $\{\mathbf{Y}^{(P_{c})}\}_{c=1}^{C}$, $\{\lambda_{c}\}_{c=1}^{C}$, $\tau$, $N_{it}$)
-2       $\mathbf{m}_{i}^{(B)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, $\mathbf{s}_{i}^{(B)}:=1$ and $\mathbf{y}_{i}^{(B)}:=\nicefrac{{1}}{{n_{c}}}$
+Our method relies on the analysis of $\theta_{B}=\{(\mathbf{m}_{i}^{(B)},\mathbf{s}_{i}^{(B)},\mathbf{v}_{i}^{(B)})\}_{i=1}^{K_{B}}\mapsto\sum_{c=1}^{C}\lambda_{c}\mathcal{SMW}_{2}(B,P_{c})^{2}$.
+First, for a fixed $\theta_{B}$, we find $\omega_{1}^{\star},\cdots,\omega_{C}^{\star}$ transport plans.
+Then, for fixed transport plans, we solve,
+$\underset{\theta_{B}}{\text{argmin}}\,\mathcal{L}(\theta_{B})$ $=\sum_{c=1}^{C}\lambda_{c}\sum_{i=1}^{K_{B}}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{\star}C_{c,i,j},$
+where $\text{where }C_{c,i,j}$ $=\lVert\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})}\rVert_{2}^{2}+\lVert\mathbf{s}_{i}^{(B)}-\mathbf{s}_{j}^{(P_{c})}\rVert_{2}^{2}+\beta\lVert\mathbf{v}_{i}^{(B)}-\mathbf{v}_{j}^{(P_{c})}\rVert_{2}^{2}$
+which can be optimized by taking derivatives with respect $\mathbf{m}^{(B)}_{i}$, $\mathbf{s}^{(B)}_{i}$ and $\mathbf{v}^{(B)}_{i}$.
+For instance, taking the derivative of $\mathcal{L}(\theta_{B})$ with respect $\mathbf{m}_{i}^{(B)}$,
+$\dfrac{\partial\mathcal{L}}{\partial\mathbf{m}_{i}^{(B)}}=2\sum_{c=1}^{C}\lambda_{k}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{\star}(\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})})=\dfrac{2}{K_{B}}\mathbf{m}_{i}^{(B)}-2\sum_{c=1}^{C}\lambda_{c}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{\star}\mathbf{m}_{j}^{(P_{c})}$
+setting the derivative to 00, one has, $\mathbf{m}_{i}^{(B)}=\sum_{c=1}^{C}\lambda_{c}T_{\omega_{c}^{\star}}(\mathbf{m}_{i}^{(B)}).$ Similar results can be acquired for $\mathbf{s}_{i}^{(B)}$ and $\mathbf{v}_{i}^{(B)}$ by taking the appropriate derivatives.
+Our strategy is shown in Algorithm  1.
+1 function smw_barycenter($\{(\mathbf{M}^{(P_{c})},\mathbf{S}^{(P_{c})},\mathbf{V}^{(P_{c})})\}_{c=1}^{C}$, $\tau$, $N_{it}$)
+2       $\mathbf{m}_{i}^{(B)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, $\mathbf{s}_{i}^{(B)}=1$ and $\mathbf{y}_{i}^{(B)}=\nicefrac{{\mathds{1}_{n_{c}}}}{{n_{c}}}$
 3       while $|L_{it}-L_{it-1}|\geq\tau$ and $it\leq N_{it}$ do
@@
             // Note: $\mathcal{W}_{2}(B_{i},P_{c,j})^{2}=\lVert\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})}\rVert_{2}^{2}+\lVert\mathbf{s}_{i}^{(B)}-\mathbf{s}_{j}^{(P_{c})}\rVert_{2}^{2}$
-7             $L_{it}=\sum_{c=1}^{C}\lambda_{c}\sum_{i=1}^{K_{B}}\sum_{j=1}^{K_{P}}\omega^{(c,it)}_{ij}\biggr{(}(\mathcal{W}_{2}(B_{i},P_{c,j})^{2}+\lVert\mathbf{y}^{(B)}_{i}-\mathbf{y}^{(P_{c})}_{j}\rVert_{2}^{2})\biggr{)}$
+7             $L_{it}=\sum_{c=1}^{C}\lambda_{c}\sum_{i=1}^{K_{B}}\sum_{j=1}^{K_{P}}\omega^{(c,it)}_{ij}\biggr{(}(\mathcal{W}_{2}(B_{i},P_{c,j})^{2}+\beta\lVert\mathbf{v}^{(B)}_{i}-\mathbf{v}^{(P_{c})}_{j}\rVert_{2}^{2})\biggr{)}$
              // Update barycenter parameters
-8             $\mathbf{m}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{k}T_{\omega^{(k,it)}}(\mathbf{m}^{(B)}_{i})$
-9             $\mathbf{s}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{k}T_{\omega^{(k,it)}}(\mathbf{s}^{(B)}_{i})$
-10             $\mathbf{y}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{k}T_{\omega^{(k,it)}}(\mathbf{y}^{(B)}_{i})$
+8             $\mathbf{m}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{c}T_{\omega^{(c,it)}}(\mathbf{m}^{(B)}_{i})$
+9             $\mathbf{s}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{c}T_{\omega^{(c,it)}}(\mathbf{s}^{(B)}_{i})$
+10             $\mathbf{v}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{c}T_{\omega^{(c,it)}}(\mathbf{v}^{(B)}_{i})$
 11
-12      return $\mathbf{M}^{(B)}$, $\mathbf{S}^{(B)}$, $\mathbf{Y}^{(B)}$
+12      return $\mathbf{M}^{(B)}$, $\mathbf{S}^{(B)}$, $\mathbf{V}^{(B)}$
 13
@@
 In this section, we detail two contributions for MSDA based on GMM-OT: GMM-WBT and GMM-DaDiL.
-In both cases, we suppose access to $N$ labeled source measures $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N}$ and an unlabeled target measure $Q_{T}$.
+In both cases, we suppose access to $N$ labeled source GMMs $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N}$ and an unlabeled target GMM $Q_{T}$.
 Contrary to the empirical versions of these algorithms [8, 9, 12], we assume that an axis-aligned GMM has been learned for each domain, including the target.
@@
 $\hat{\mathbf{m}}_{i}^{(Q_{T})}=K_{B}\sum_{j=1}^{K_{T}}\omega_{ij}^{(T)}\mathbf{m}_{j}^{(Q_{T})}\text{, and, }\hat{\mathbf{s}}_{i}^{(Q_{T})}=K_{B}\sum_{j=1}^{K_{T}}\omega_{ij}^{(T)}\mathbf{s}_{j}^{(Q_{T})}.$ (15)
-With a labeled GMM, $\{\hat{\mathbf{m}}_{i}^{(Q_{T})},\hat{\mathbf{s}}_{i}^{(Q_{T})},\mathbf{y}_{i}^{(B)}\}_{i=1}^{K_{B}}$, on the target domain, we can learn a classifier on the target domain as explained in section 3.2.
-1 function gmm_dadil($\{\mathbf{M}^{(Q_{S_{\ell}})}\}_{\ell=1}^{N}$, $\{\mathbf{S}^{(Q_{S_{\ell}})}\}_{\ell=1}^{N}$, $\{\mathbf{Y}^{(Q_{S_{\ell}})}\}_{\ell=1}^{N}$, $N_{it}$, $\eta$)
+With a labeled GMM, $\{\hat{\mathbf{m}}_{i}^{(Q_{T})},\hat{\mathbf{s}}_{i}^{(Q_{T})},\mathbf{v}_{i}^{(B)}\}_{i=1}^{K_{B}}$, on the target domain, we can learn a classifier on the target domain as explained in section 3.2.
+1 function gmm_dadil($\{(\mathbf{M}^{(Q_{S_{\ell}})},\mathbf{S}^{(Q_{S_{\ell}})},\mathbf{V}^{(Q_{S_{\ell}})})\}_{\ell=1}^{N}$, $\{(\mathbf{M}^{(Q_{T})},\mathbf{S}^{(Q_{T})})\}$, $N_{it}$, $\eta$)
        // Initialization.
@@
 3       for $it=1,\cdots,N_{it}$ do
-4             $L=0$
+4             $L\leftarrow 0$
+             // Change of variables
+5             $\mathbf{v}_{i}^{(P_{k})}\leftarrow\text{softmax}(\mathbf{u}_{i}^{(P_{k})})$
              // Evaluate supervised loss on sources
@@
 7                   $L\leftarrow L+\mathcal{SMW}_{2}(Q_{S_{\ell}},\mathcal{B}(\lambda_{\ell},\mathcal{P}))^{2}$
-7
+8
             // Evaluate unsupervised loss on targets
@@
              // Gradient step
-9             $\mathbf{x}_{j}^{(P_{k})}\leftarrow\mathbf{x}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{x}^{(P_{k})}_{j}}}$
-10             $\mathbf{p}_{j}^{(P_{k})}\leftarrow\mathbf{u}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{u}^{(P_{k})}_{j}}}$
+10             $\mathbf{m}_{j}^{(P_{k})}\leftarrow\mathbf{m}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{m}^{(P_{k})}_{j}}}$
+11             $\mathbf{u}_{j}^{(P_{k})}\leftarrow\mathbf{u}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{u}^{(P_{k})}_{j}}}$
              // Note: we project variables $\mathbf{s}$ and $\lambda$.
 12             $\mathbf{s}_{j}^{(P_{k})}\leftarrow\text{proj}_{\mathbb{R}^{d}_{+}}(\mathbf{s}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{s}^{(P_{k})}_{j}}})$
-12             $\lambda_{\ell}\leftarrow\text{proj}_{\Delta_{K}}(\lambda_{\ell}-\eta\nicefrac{{\partial L}}{{\partial\lambda_{\ell}}})$
-13
-14      return $\mathcal{P},\Lambda$
-15
+13             $\lambda_{\ell}\leftarrow\text{proj}_{\Delta_{C}}(\lambda_{\ell}-\eta\nicefrac{{\partial L}}{{\partial\lambda_{\ell}}})$
+14
+15      return $\Lambda,\mathcal{P}$
+16
 Algorithm 2 GMM-Dataset Dictionary Learning
@@
 Our second algorithm consists of a parametric version for the DaDiL algorithm of [12].
-The idea is to replace the atoms in $\mathcal{P}=\{\hat{P}_{c}\}_{c=1}^{C}$ by GMMs parametrized through $\Theta_{P}=\{(\mathbf{M}^{(P_{c})},\mathbf{S}^{(P_{c})},\mathbf{Y}^{(P_{c})})\}_{c=1}^{C}$.
+The idea is to replace the atoms in $\mathcal{P}=\{\hat{P}_{c}\}_{c=1}^{C}$ by GMMs parametrized through $\Theta_{P}=\{(\mathbf{M}^{(P_{c})},\mathbf{S}^{(P_{c})},\mathbf{V}^{(P_{c})})\}_{c=1}^{C}$.
 Learning a dictionary is thus equivalent to estimating these parameters, that is,
-$\mathcal{L}(\Lambda,\Theta_{P})$ $=\underset{\Lambda,\mathcal{P}}{\text{argmin}}\,\mathcal{MW}_{2}(Q_{T},\mathcal{B}(\lambda_{T};\mathcal{P}))^{2}+\sum_{\ell=1}^{N}\mathcal{SMW}_{2}(Q_{\ell},\mathcal{B}(\lambda_{\ell};\mathcal{P}))^{2}.$ (16)
-An important feature of this strategy is that the barycenters $\mathcal{B}$ inherit the labels from the atoms $\mathcal{P}$.
-As a result, we can estimate a labeled GMM for $Q_{T}$ in one step through $\mathcal{B}(\lambda_{T};\mathcal{P})$.
+$(\Lambda^{\star},\Theta_{P}^{\star})$ $=\underset{\Lambda,\Theta_{P}}{\text{argmin}}\,\mathcal{MW}_{2}(Q_{T},\mathcal{B}(\lambda_{T},\mathcal{P}))^{2}+\sum_{\ell=1}^{N}\mathcal{SMW}_{2}(Q_{\ell},\mathcal{B}(\lambda_{\ell};\mathcal{P}))^{2}.$ (16)
 While eq.
@@
 Once the dictionary $(\Lambda,\mathcal{P})$ is learned, we are able to reconstruct the domains in MSDA via the barycenter $\mathcal{B}(\lambda;\mathcal{P})$.
-We are especially interested in the target reconstruction $\lambda_{T}$, i.e., $\mathcal{B}(\lambda_{T};\mathcal{P})$.
+We are especially interested in the target reconstruction $\lambda_{T}$, i.e., $\mathcal{B}(\lambda_{T},\mathcal{P})$.
 This barycenter is a labeled GMM (as we show in figure 4b).
 As a result, we can obtain labeled samples from this GMM, then use them to train a classifier that works on the target domain.
-Faster Dataset Dictionary Learning.
 The computational complexity of an optimization step of algorithm 2 corresponds to $\mathcal{O}(N\times N_{it}\times C\times K^{3}\log K)$, i.e., we calculate $N$ barycenters of $C$ atoms.
@@
 In our experiments in section 4, we show that we achieve state-of-the-art performance with $K$ on the same order of magnitude as $n_{b}$ (e.g., a few hundred Gaussian components).
-As a result, we achieve a speed-up on the order of $M$ while solving an exact OT problem.
+As a result, we achieve a speed-up on the order of $M$ while solving an exact OT problem (see Figure 5 (c) below).
 4 Experiments
-In this section, we present our experiments.
-First, we present a toy example, where the source and target measures were generated by transforming an initial measure.
-As such, this toy example illustrates our goal to model distributional shift.
-Second, we experiment with benchmarks in MSDA.
-Figure 3: Data used in the toy experiment. Each of these datasets was generated by applying an affine transformation to an initial dataset.
-Toy Example.
-In figure 3 we show the generated datasets.
-Starting with GMM-WBT, figure 4(a) shows the learned GMMs for each dataset, in which the target GMM is not labeled.
-The barycenter of $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N}$ is shown in figure 4(b) (bottom-left).
+4.1 Toy Example
+In this section, we explore GMM-WBT and GMM-DaDiL in the context of a toy example.
+We generate 4 datasets over $\mathbb{R}^{d}$, by gradually shifting and deforming initial measure through an affine mapping.
+In figure 3 (a) we show the generated datasets.
+Starting with GMM-WBT, figure 3 (b) shows the learned GMMs for each dataset, in which the target GMM is not labeled.
+The barycenter of $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N}$ is shown in figure 3(c) (bottom-left).
 This barycenter is labeled.
 As a result, we may transfer its parameters to the target domain through GMM-OT (upper part), which leads to a labeled GMM in the target domain.
-(a) GMMs learned on each domain.
-(b) GMM-WBT.
-Figure 4: Summary of GMM-WBT, in which a labeled GMM is determined on the target domain by transporting the parameters of the barycenter of sources.
+(a) Data.
+(b) GMMs.
+(c) GMM-WBT
+Figure 3: Data and GMMs used in the toy experiment.
+In (a) Each of these datasets was generated by applying an affine transformation to an initial dataset.
+In (b), we show an axis-aligned GMM fitted to the data via EM.
+In (c), we show a summary of GMM-WBT, where show the OT plan between components (upper part) between $B$ (left) and $Q_{T}$ (right).
+The resulting labeled GMM is shown in the lower right part of (c).
 Next, we show in figure 4 a summary for the GMM-DaDiL optimization process (figure 4(a)), and the reconstruction of target domain GMMs (figure 4(b)).
@@
 (b) Reconstructions ($it=200$).
-Figure 5: Optimization of GMM-DaDiL and reconstruction of domain GMMs.
-Experimental Setting.
+Figure 4: Optimization and reconstruction summaries using GMM-DaDiL.
+In (a), we show the evolution of loss, negative log-likelihood and barycentric coordinates (i.e., $\lambda_{\ell}$) over the course of optimization.
+In (b), we show the reconstructed GMMs (i.e., $\mathcal{B}(\lambda_{\ell},\mathcal{P})$) when the algorithm converges.
+4.2 Multi-Source Domain Adaptation
 We compare our method to prior art.
@@
 LTC-MSDA 82.21 $\pm$ 8.03 75.33 $\pm$ 5.91 81.04 $\pm$ 5.45 79.52
-KD3A 81.02 $\pm$ 2.92 78.04 $\pm$ 4.05 74.64 $\pm$ 5.65 77.90
+KD3A§ 81.02 $\pm$ 2.92 78.04 $\pm$ 4.05 74.64 $\pm$ 5.65 77.90
 Co-MDA 62.66 $\pm$ 0.96 55.78 $\pm$ 0.85 76.35 $\pm$ 0.79 64.93
@@
 M3SDA 81.17 $\pm$ 2.00 61.61 $\pm$ 2.71 79.99 $\pm$ 2.71 79.12 $\pm$ 2.41 75.16 $\pm$ 3.01 78.91 $\pm$ 3.24 75.99
-KD3A 72.52 $\pm$ 3.04 18.96 $\pm$ 4.54 81.02 $\pm$ 2.40 74.42 $\pm$ 1.60 67.18 $\pm$ 2.37 78.22 $\pm$ 2.14 65.38
+KD3A§ 72.52 $\pm$ 3.04 18.96 $\pm$ 4.54 81.02 $\pm$ 2.40 74.42 $\pm$ 1.60 67.18 $\pm$ 2.37 78.22 $\pm$ 2.14 65.38
 Co-MDA 64.56 $\pm$ 0.62 35.99 $\pm$ 1.21 79.66 $\pm$ 1.36 72.06 $\pm$ 1.66 66.33 $\pm$ 0.97 78.91 $\pm$ 1.87 66.34
@@
 (d) TEP.
-Benchmarking Results.
 First, OT-based methods generally outperform other methods in MSDA.
-Overall, these methods solve a simpler task compared to deep DA methods, as they do not need to update the encoder network during adaptation.
+Overall, shallow DA methods solve a simpler task compared to deep DA methods, as they do not need to update the encoder network during adaptation.
 Second, the GMM-OT framework generally improves over using empirical OT.
@@
 Furthermore, GMM-DaDiL outperforms its empirical counterpart on all benchmarks, as well as GMM-WBT.
-This point further showcases the power of dictionary learning in MSDA.
-Figure 6: In (a), we analyse the performance of interpolations $\mathcal{B}((\lambda_{0},1-\lambda_{0});\mathcal{Q}_{S})$, $\mathcal{Q}_{S}=\{Q_{S_{1}},Q_{S_{2}}\}$ and $\mathcal{B}((\lambda_{0},1-\lambda_{0});\mathcal{P})$ with learned $\mathcal{P}=\{P_{1},P_{2}\}$ for GMM-WBT and GMM-DaDiL.
+This point further illustrates the power of dictionary learning in MSDA.
+Note that, in table 2(d) (d), GMM-DaDiL manages to have the best average adaptation performance across domains without actually being the best on any single domain.
+As a consequence, GMM-DaDiL enjoys better stability, with respect distribution shift, than previous methods.
+4.3 Lighter, Better, Faster Domain Adaptation
+(a) Lighter.
+(b) Better.
+(c) Faster.
+Figure 5: Lighter, Better, Faster.
+In (a), we analyse the performance of interpolations $\mathcal{B}((\lambda_{0},1-\lambda_{0});\mathcal{Q}_{S})$, $\mathcal{Q}_{S}=\{Q_{S_{1}},Q_{S_{2}}\}$ and $\mathcal{B}((\lambda_{0},1-\lambda_{0});\mathcal{P})$ with learned $\mathcal{P}=\{P_{1},P_{2}\}$ for GMM-WBT and GMM-DaDiL.
 In (b), we analyse the efficiency of barycenter-based methods under an increasing number of GMM components (number of samples for DaDiL and WBT).
 GMM-DaDiL has state-of-the-art performance even for the extreme case where $K=65$.
-Better MSDA.
-We design an experiment to further showcase the contribution of dictionary learning for MSDA.
-We use the adaptation $(D,W)\rightarrow A$ in the Office-31 benchmark as the basis of our experiment.
-Note that GMM-DaDiL reconstructs the target domain via a barycenter $\mathcal{B}(\lambda_{T};\mathcal{P})$, where $\lambda_{T}$ and $\mathcal{P}$ are learned parameters.
-We thus ablate the learning of $\lambda_{T}$ and $\mathcal{P}$, i.e., we compare it to $\lambda_{T}=(\lambda_{0},1-\lambda_{0})$, $\lambda_{0}\in[0,1]$ and $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N_{S}}$.
-This generates a series of measures parametrized by $\lambda_{0}$.
-To further match $\mathcal{B}(\lambda_{T};\mathcal{Q}_{S})$ with $Q_{T}$, we transport it to $Q_{T}$ through eq.
-15.
-Note that this corresponds to performing GMM-WBT with a barycenter calculated with $\lambda_{T}$.
-The overall experiment is shown in figure 6.
-While the performance of GMM-WBT remains approximately stable, that of GMM-DaDiL grows as we move closer to $\lambda_{T}^{\star}$ learned by dictionary learning (blue star).
-Overall, the interpolation space generated by atoms better captures the distributional shift occurring on the target domain.
-Lighter MSDA.
-We further investigate the adaptation efficiency of our methods with respect the number of GMM components $n$.
-We use the adaptation task $(Cl,Pr,Rw)\rightarrow Ar$ from Office-Home to showcase our analysis.
-Note that this parameter is analogous to the number of samples in DaDiL or WBT.
-We thus show a comparison in figure 6 b.
-From this figure, we see that GMM-DaDiL surpasses all other methods over the entire range $n\in\{65,130,\cdots,910\}$, which showcases the advantage of using GMMs and dictionary learning.
-Especially, its empirical counterpart, DaDiL, needs a large number of samples in order to accurately represent probability measures.
+In (c), we compare the running time of GMM-DaDiL with that of DaDiL, as a function of number of components $K$ and batch size $n_{b}$, respectively.
+This figure illustrates the speedup of GMM-DaDiL as the number of samples in DaDiL (and hence, $M=\lceil\nicefrac{{n}}{{n_{b}}}\rceil$) increases.
+Circles represent the average over $5$ independent runs, while the error bars show $2$ times the standard deviation.
+Our first experiment illustrates why GMM-DaDiL is lighter than previous barycenter-based algorithms, such as DaDiL.
+In this context, a lighter model needs less parameters to achieve a certain domain adaptation performance.
+We rank GMM-OT models by the number of components $K$, and empirical models by the number of samples $n$ in their support.
+Note that these parameters regulate the complexity of these algorithms.
+We use the adaptation task $(Cl,Pr$,$Rw)\rightarrow Ar$ from Office-Home for our analysis.
+We show a comparison in figure 5 (a).
+From this figure, we see that GMM-DaDiL surpasses all other methods over the entire range $K\in\{65,130,\cdots,910\}$.
+Especially, its empirical counterpart, DaDiL, needs a large number of samples for accurately represent probability measures.
 Curiously, the performance of GMM-WBT and WBT are quite similar.
@@
 As a result, in this adaptation task, the GMM version of WBT has similar performance to the empirical version.
+Our second experiment illustrates why GMM-Optimal Transport Domain Adaptation (OTDA) provides a better framework for MSDA.
+We use the adaptation $(D,W)\rightarrow A$ in the Office-31 benchmark as the basis of our experiment.
+Note that GMM-DaDiL reconstructs the target domain via a barycenter $\mathcal{B}(\lambda_{T},\mathcal{P})$, where $\lambda_{T}$ and $\mathcal{P}$ are learned parameters.
+We thus ablate the learning of $\lambda_{T}$ and $\mathcal{P}$, i.e., we compare it to $\lambda_{T}=(\lambda_{0},1-\lambda_{0})$, $\lambda_{0}\in[0,1]$ and $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N_{S}}$.
+This generates a series of measures parametrized by $\lambda_{0}$.
+To further match $\mathcal{B}(\lambda_{T},\mathcal{Q}_{S})$ with $Q_{T}$, we transport it to $Q_{T}$ through eq.
+15.
+Note that this corresponds to performing GMM-WBT with a barycenter calculated with $\lambda_{T}$.
+The overall experiment is shown in figure 5 (b).
+While the performance of GMM-WBT remains approximately stable, that of GMM-DaDiL grows as we move closer to $\lambda_{T}^{\star}$ learned by dictionary learning (blue star).
+Overall, the interpolation space generated by atoms better captures the distributional shift occurring on the target domain.
+Our third experiment shows that GMM-DaDiL is faster than DaDiL.
+We plot the running time of these methods on the Office 31 benchmark, for the $(D,W)\rightarrow A$ adaptation task.
+The variables that influence the complexity of GMM and empirical DaDiL are the number of components $K$ and the batch size $n_{b}$, respectively.
+For GMM-DaDiL, we simply measure its running time for $5$ independent runs of the algorithm (blue curve) for each $K\in\{31,62,\cdots,217\}$.
+For DaDiL, we set $n_{b}\in\{31,62,\cdots,217\}$, and set $n=M\times n_{b}$, where $M$ is the number of mini-batches.
+We measure the performance over $5$ independent runs as well.
+Other than these parameters, we fix $N_{iter}=50$ and $C=3$.
+As shown in Figure 5 (c), the running time of GMM-DaDiL and DaDiL are essentially equivalent for $M=1$.
+For $M>1$, we have a speedup that is proportional to $M$.
+In our fourth experiment, we use the $(B,C)\rightarrow A$ adaptation task of CWRU.
+We are interested in visualizing the evolution of atoms and reconstructions with respect DaDiL and GMM-DaDiL iterations.
+We visualize this evolution through UMAP [33], i.e., we concatenate the data from DaDiL’s atoms, i.e., $\mathbf{x}_{i}^{(P_{c,it})}$, so that these are jointly embedded into $\mathbb{R}^{2}$.
+For GMM-DaDiL, we concatenate the mean parameters, i.e., $\mathbf{m}_{i}^{(P_{c,it})}$.
+We summarize our results in figure 6.
+Overall, as shown in figure 6 (a – d), GMM-DaDiL optimization is more stable than that of DaDiL, especially since we do not use mini-batches.
+This remark is also evidenced in the reconstructions in figure 6 (e – f).
 (a) $P_{1,it}$
@@
 In (g, h), we show the reconstructions alongside target domain data at convergence.
-Visualization.
-In this experiment, we use the $(B,C)\rightarrow A$ adaptation task of CWRU.
-We are interested in visualizing the evolution of atoms and reconstructions with respect DaDiL and GMM-DaDiL iterations.
-We visualize this evolution through UMAP [32], i.e., we concatenate the data from DaDiL’s atoms, i.e., $\mathbf{x}_{i}^{(P_{c,it})}$, so that these are jointly embedded into $\mathbb{R}^{2}$.
-For GMM-DaDiL, we concatenate the mean parameters, i.e., $\mathbf{m}_{i}^{(P_{c,it})}$.
-We summarize our results in figure 7.
-Overall, as shown in figure 7 (a – d), GMM-DaDiL optimization is more stable than that of DaDiL, especially since we do not use mini-batches.
-This remark is also evidenced in the reconstructions in figure 7 (e – f).
 5 Conclusion
@@
 Overall, we propose methods that are lighter, better, faster than previous empirical OT methods in MSDA.
-Appendix A Introduction
-In this supplementary materials, we provide additional details about our theoretical and experimental results.
-Appendix B Proofs
-B.1 Proof for Theorem 1
-Theorem 3.
-Let $P$ and $Q$ be two GMMs with components $P_{i}=\mathcal{N}(\mathbf{m}_{i}^{(P)},(\mathbf{s}_{i}^{(P)})^{2})$ (resp.
-$Q_{j}$) and $\omega^{\star}$ be the solution of eq.
-10.
-The first-order optimality conditions of $\mathcal{MW}_{2}^{2}$, with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by,
-$\mathbf{m}_{i}=T_{\omega}(\mathbf{m}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{m}_{j}^{(Q)}\text{, and }\mathbf{s}_{i}=T_{\omega}(\mathbf{s}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{s}_{j}^{(Q)}.$ (17)
-Proof.
-Under the GMM-OT framework, the $2-$mixture Wasserstein distance is given by,
-$\mathcal{MW}_{2}(P,Q)^{2}=\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathcal{W}_{2}(P_{i},Q_{j})^{2},$
-where $\omega_{ij}^{\star}$ is the mass transported from component $i$ in $P$ to component $j$ in $Q$ by the solution to the GMM-OT problem.
-Since $P_{i}$ and $Q_{j}$ are axis-aligned Gaussian measures,
-$\mathcal{W}_{2}(P_{i},Q_{j})^{2}=\lVert\mathbf{m}_{i}-\mathbf{m}_{j}^{(Q)}\rVert_{2}^{2}+\lVert\mathbf{s}_{i}-\mathbf{s}_{j}^{(Q)}\rVert_{2}^{2},$
-where $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are the free parameters we want to determine.
-Henceforth we derive the first-order conditions for $\mathbf{m}_{i}$.
-The derivations for $\mathbf{s}_{i}$ are completely analogous.
-By linearity,
-$\dfrac{\partial\mathcal{MW}_{2}^{2}}{\partial\mathbf{m}_{i}}=2\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}(\mathbf{m}_{i}-\mathbf{m}_{j}^{(Q)})=2\biggr{(}p_{i}\mathbf{m}_{i}-\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathbf{m}_{j}^{(Q)}\biggr{)},$
-equating this last term to 00, one gets the desired equality. ∎
-The result of theorem 3 is quite similar to the barycentric mapping of [33, 6] which was previously used to map points in $P$ to points in $Q$. Here, we transport the parameters of GMMs instead.
-B.2 Proof of Theorem 2
-Theorem 4.
-Under the same conditions of theorem 3, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{y}_{i}^{(P)}$ and $\mathbf{y}_{j}^{(Q)}$.
-The first order optimality conditions of $\mathcal{SMW}_{2}$ with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by eq.
-17.
-Furthermore, for $\mathbf{y}_{i}$,
-$\mathbf{y}_{i}=T_{\omega}(\mathbf{y}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{y}_{j}^{(Q)}.$ (18)
-Proof.
-Note that the label distance term added to $\mathcal{SMW}$ does not interfere with the terms on $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$, hence the optimality conditions with respect these variables remain unchanged.
-Due the independence of these terms, the first-order optimality condition with respect $\mathbf{y}_{i}$ is,
-$\dfrac{\partial\mathcal{SMW}_{2}^{2}}{\partial\mathbf{y}_{i}}=2\beta\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}(\mathbf{y}_{i}-\mathbf{y}_{j}^{(Q)})=2\beta\biggr{(}p_{i}\mathbf{y}_{i}-\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathbf{y}_{j}^{(Q)}\biggr{)},$
-which, for $\beta>0$, is zero if and only if $\mathbf{y}_{i}=T_{\omega}(\mathbf{y}_{i}^{(P)})$. ∎
-B.3 Derivation of Algorithm 1
-In this section, we derive the first-order conditions for the functional $B\mapsto\sum_{c}\lambda_{c}\mathcal{SMW}_{2}(B,P_{c})^{2}$.
-Let $B$ be parametrized by $\theta_{B}=\{(\mathbf{m}_{i}^{(B)},\mathbf{s}_{i}^{(B)},\mathbf{y}_{i}^{(B)})\}_{i=1}^{K_{B}}$, where these variables are initialized randomly as in algorithm 1.
-For fixed $\theta_{B}$, we find $\omega_{1}^{\star},\cdots,\omega_{C}^{\star}$ transport plans.
-Then, for fixed transport plans, we solve,
-$\underset{\theta_{B}}{\text{argmin}}\,\mathcal{L}(\theta_{B})$ $=\sum_{c=1}^{C}\lambda_{c}\sum_{i=1}^{K_{B}}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{(\star)}M_{c,i,j},$
-where $\text{where }M_{c,i,j}$ $=\lVert\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})}\rVert_{2}^{2}+\lVert\mathbf{s}_{i}^{(B)}-\mathbf{s}_{j}^{(P_{c})}\rVert_{2}^{2}+\beta\lVert\mathbf{y}_{i}^{(B)}-\mathbf{y}_{j}^{(P_{c})}\rVert_{2}^{2}$
-which can be optimized by taking derivatives with respect $\mathbf{m}^{(B)}_{i}$, $\mathbf{s}^{(B)}_{i}$ and $\mathbf{y}^{(B)}_{i}$.
-For instance, taking the derivative of $\mathcal{L}(\theta_{B})$ with respect $\mathbf{m}_{i}^{(B)}$,
-$\dfrac{\partial\mathcal{L}}{\partial\mathbf{m}_{i}^{(B)}}=2\sum_{c=1}^{C}\lambda_{k}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{\star}(\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})})=\dfrac{2}{K_{B}}\mathbf{m}_{i}^{(B)}-2\sum_{c=1}^{C}\lambda_{c}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{\star}\mathbf{m}_{j}^{(P_{c})}$
-setting the derivative to 00, one has,
-$\mathbf{m}_{i}^{(B)}$ $=\sum_{c=1}^{C}\lambda_{c}T_{\omega_{c}^{\star}}(\mathbf{m}_{i}^{(B)}).$
-Similar results can be acquired for $\mathbf{s}_{i}^{(B)}$ and $\mathbf{y}_{i}^{(B)}$ by taking the appropriate derivatives.
-Appendix C Additional Experiments
-Figure 8: Evolution of barycentric coordinates over the simplex $\Delta_{C}$.
-Toy Example.
-We analyze the evolution of the atoms in GMM-DaDiL as training progresses (figure 9).
-Note that, initially, the means are close to each other, so there are no distinction between classes.
-Likewise, in our initialization, the labels are uniform histograms over the classes, so that the atoms are not certain about which class they represent.
-With 50 iterations, DaDiL manages to spread the class clusters, but, at this point, the barycentric coordinates of each domain are still too similar (i.e., $\lambda_{\ell}\approx\mathds{1}_{K}/K$, see figure 2.a in the main paper).
-As the training progresses, the barycentric coordinates of the different domains begins to spread over the simplex $\Delta_{C}$ (see figure 8).
-As a result, the atoms begin to be different from each other (compare figures 9 (b) and (c)).
-Hyper-parameter Setting: Here we detail the hyper-parameter setting for our method.
-GMM-WBT involves a single hyper-parameter: the number of components $K$ in the GMMs.
-DaDiL involves a few other parameters: the number of atoms $C$, the learning rate $\eta$ and the number of optimization steps $N_{iter}$.
-As [12], we use Adam [34] for optimizing eq.
-16 (see main paper).
-Over all our experiments we set $\eta=10^{-1}$.
-We show our hyper-parameters in table 3.
-We remark that, as we analyze in Fig 6.
-b (main paper), our proposed methods, especially GMM-DaDiL, are robust to the choice of hyper-parameters.
-Table 3: Hyper-parameter setting
-Benchmark $K$ (GMM-WBT) $K$ (GMM-DaDiL) $C$ $N_{iter}$
-Office 31 155 186 3 150
-Office-Home 910 455 4 150
-TEP 261 174 7 150
-CWRU 30 90 5 200
-(a) 1st iteration.
-(b) 50th iteration.
-(c) 100th iteration.
-(d) 150th iteration.
-(e) 200th iteration.
-Figure 9: Evolution of atoms as the training progresses.